### v1

# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

Abstract.
Advances in eye-tracking control for assistive robotic arms provide intuitive interaction opportunities for people with physical disabilities. Shared control has gained interest in recent years by improving user satisfaction through partial automation of robot control. We present an eye-tracking-guided shared control design based on insights from state-of-the-art literature. A Wizard of Oz setup was used in which automation was simulated by an experimenter to evaluate the concept without requiring full implementation. This approach allowed for rapid exploration of user needs and expectations to inform future iterations. Two studies were conducted to assess user experience, identify design challenges, and find improvements to ensure usability and accessibility. The first study involved people with disabilities by providing a survey, and the second study used the Wizard of Oz design in person to gain technical insights, leading to a comprehensive picture of findings.
Assistive Robotics, Eye-Tracking, Robot Manipulation, Assistive Device, User Experience, Disability, Wizard of Oz
ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Pointing devicesccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Development frameworks and environments
1. Introduction
People with physical disabilities such as locked-in syndrome, cerebral palsy, late-stage amyotrophic lateral sclerosis, or multiple sclerosis benefit from non-haptic controls such as Brain-Computer Interfaces and eye-tracking controllers. Implementation as control inputs for assistive robotic arms can enhance user independence, self-determination, and privacy in daily life by reducing the dependency on family members and caregivers (Fischer-Janzen et al., 2024; Awuah et al., 2024; Baumeister et al., 2021).
Regardless of the controlled technology, challenges have been identified with such input devices in the areas of safety, availability of participants for clinical trials, privacy, interdisciplinarity, and cost (Maiseli et al., 2023; Suryadarma et al., 2024; Fischer-Janzen et al., 2024; Nanavati et al., 2024b). This has led to the development of a variety of robot control methods that differ in terms of input modalities and control strategies (Mohebbi, 2020; Fischer-Janzen et al., 2024). Depending on the input modality, the achievable varying level of robot autonomy yields different advantages and disadvantages. For instance, direct control of the robot leads to the need for mode switches to cover the 3-dimensional space the robot can move, or to control all the joints of the robot, enhancing the needed concentration on the tasks (Fischer-Janzen et al., 2024; Kronhardt et al., 2023; Herlant et al., 2016). The implementation of a shared control approach, in which tasks are partially automated, has been demonstrated to effectively address this issue. Such approaches have been shown to positively impact interaction perception and reduce the time required to complete tasks (Bhattacharjee et al., 2020; Kronhardt et al., 2023). A potential challenge associated with a high level of automation is user discomfort, which can arise if users feel that they are losing control over the system. In this context, community feedback is essential to improve design and evaluate usability and accessibility (Stalljann et al., 2020; Baumeister et al., 2021; Nanavati et al., 2024a). It provides firsthand insights into user expectations and perceptions of their daily lives, thus enhancing acceptance rates (Nanavati et al., 2024a; Canal et al., 2021).
This work incorporates feedback from the community on an initial shared control design to identify challenges and improve the design. A Wizard of Oz design of the eye-tracking-driven control was developed, with which two studies were conducted. The first study was an online survey distributed to people with disabilities, their family members and friends, and caregivers. The objective was to assess the usability in daily life and to identify the tasks that participants would prefer to perform. The stakeholder group was defined as individuals with physical disabilities who are unable to use their extremities to control the robot. In the second study, non-disabled participants with varying levels of robotics experience interacted with the system to obtain technical parameters that would provide constraints for a future control algorithm. In this context technical parameters describe for example the maximum and minimum dwell time and the level of accuracy needed for a robust control. The main contributions of this work are as follows:
  1. (1)
    Presentation of a shared control design for everyday tasks.
  2. (2)
    An evaluation of functionality and usability through opinions of potential users, their families, caregivers and non-disabled participants.
  3. (3)
    Improvements in the initial shared control design gained from the studies.
2. Background
2.1. Shared control for robotic arms
”Shared control in assistive robotics blends human autonomy with computer assistance, thus simplifying complex tasks for individuals with physical impairments” (Goldau et al., 2024). Advances in sensing, inference, modeling, and learning methods have lead to novel developments in the research area (Selvaggio et al., 2021). In this context, researchers are investigating impact factors on shared control. Literature has been reviewed to find such factors influencing human-robot-interaction. three categories were introduced to facilitate the overview over the found outcomes: Field of use, technical challenges, level of autonomy, and user individuality and preferences.
The field of use establishes the boundaries within which shared controls are to be utilized. Robots can be used for daily tasks in unstructured environments (Stanger et al., 1994), household chores (Herr et al., 2016), and personal tasks such as personal hygiene or medical procedures (Wang et al., 2020). This determines the positioning of the robot, whether it must be attached to an electric wheelchair, on the bedside, or on a table (Stanger et al., 1994), and leads to the initial setup configuration and limitation in robot choice.
The technical challenges include reliability and functionality of the system: Safety, robustness and success rate in e.g. task completion is a standard that has to be ensured (Stanger et al., 1994; Wang et al., 2020) and the requirements are defined in ISO 13482 (für Normung e.V., DIN). The system must be robust to changing or unorganized environments, especially when the robot is mounted on an electric wheelchair (Selvaggio et al., 2021; Stanger et al., 1994). The system should be easy to learn and use, which also depends on the complexity of the task (Stanger et al., 1994; Shafti et al., 2019; Goldau et al., 2024). Visualization with an interface can help (Shafti et al., 2019; Nanavati et al., 2024b). If interfaces in the form of displays are not available other feedback methods can ensure the ease of use by communicating information to the user and reducing uncertainty in using the robot (Shafti et al., 2019; Nanavati et al., 2024b). Autonomous robotic task learning, for example imitation learning, ensures the availability of a large range of tasks (Stanger et al., 1994; Selvaggio et al., 2021). The availability of general tasks adapted to the field of usage eases the distress in the lack of capabilities (Brose et al., 2010). The appearance and aesthetics can influence the results of user’s system perception. In the past, connections were found that due to insufficient aesthetics, users expected reduced social communication and isolation (Stanger et al., 1994; Brose et al., 2010).
Ongoing debates revolve around the level of autonomy and can vary from teleoperated systems to fully autonomous robots (Beer et al., 2014). Recent studies have shown that more autonomy is not always better for the user (Nanavati et al., 2024b; Bhattacharjee et al., 2020). Autonomous behavior of the robot ”allows reducing the human operator’s workload when performing a task that can be repetitive and/or requires skills, effort or precision levels that exceed those of a human” (Selvaggio et al., 2021). If the robot produces more errors in task execution with a high level of autonomy, users tend to reject the system (Bhattacharjee et al., 2020). It can further lead to lost trust or a sense of lost control with unpredictable behavior of the system (Nanavati et al., 2024b). Validation functions communicating and verifying the next steps with the user can ease this issue (Bhattacharjee et al., 2020). On the other hand, the user’s attitude towards a task influences the wish for a higher level of autonomy or a complete delegation of the task to the robot, e.g. if the user chooses not to perform the task (Herr et al., 2016). In eye-tracking controls these levels of autonomy can be distinguished in teleoperated system, directionally controlled systems and systems with automated trajectory or grasp planning. This variation strongly influences the design process (Fischer-Janzen et al., 2024). Directional gaze control is easy to learn, provides a sense of control, and reduces cognitive load (Dragomir et al., 2021; Dziemian et al., 2016). On the other hand, it is more time consuming to control each of the robot movements separately. Shared controls are mostly controlled by the user selecting the object via fixation, indicating the wish to interact with (Catalán et al., 2017; Zeng et al., 2017; Tostado et al., 2016).
The user’s individuality and preferences vary with personal metrics such as the user’s age, level of disability, living arrangements, and familiarity with technology and affect the system’s performance perception (Stanger et al., 1994). The adaptation of the system to user preferences is mentioned by various authors (Bhattacharjee et al., 2020; Selvaggio et al., 2021). The ability to adapt the system gave rise to Shared Autonomy approaches, ”where the robot is capable of seamless adaptation of its autonomy level based on its own understanding of the human actions/intentions and of the surrounding environment” (Selvaggio et al., 2021). Incorporating the preferences of the user influences the perception and user satisfaction of the system (Cakmak et al., 2011; Canal et al., 2021). An example is the participant’s preference for handing an object with a particular orientation (Cakmak et al., 2011). A common factor, as known from the use of PCs, is the adaptation of the system to the user’s preferences by offering a settings menu. Parameters such as the velocity of the robot should be adaptable (Bhattacharjee et al., 2020). Another factor is that offering a variation of input modalities (e.g. head or eye-tracking control instead of haptic control) makes it possible to tailor the system to the user’s needs (Bhattacharjee et al., 2020; Nanavati et al., 2024b). Users want their interactions with the robot ”to be customized to their impairment, their preferences, and their context” (Nanavati et al., 2024b).
Other impact factors are a low maintenance system to ensure a reduced workload for others, such as family or caregivers (Bhattacharjee et al., 2020). Related to market introduction are factors such as the initial cost and the maintenance cost of the system (Stanger et al., 1994; Brose et al., 2010). These may have less impact on the outcomes of the user survey, but are an important factor in the design of the system.
Due to the closeness between human and robot these technical and ethical challenges coexist (Canal et al., 2021). In designing a well-balanced shared control, a trade-off must be found for all factors relating to technical and ethical challenges. In the following, we will present a solution that considers and balances the different influences.
2.2. Task preferences in assistive robot manipulation
Among the many daily tasks, research has identified the areas where people with physical disabilities prefer assistance. This can have significant impact on the design of the system, through the adaptation of grippers or, in the case of shared control, task automation.
The four most frequently mentioned tasks that an assistive robot should perform, were food related tasks (serving a meal, eating) (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Canal et al., 2021; Bhattacharjee et al., 2020; Nanavati et al., 2024a), drinking related tasks (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003), picking up objects (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Nanavati et al., 2024b), and personal hygiene (Stanger et al., 1994; Brose et al., 2010; Nanavati et al., 2024b). Picking up an object was further related to the location or a manipulation that was desired with the object such as picking it up from the floor or from a shelf, bringing it near, or moving it. Personal hygiene included various tasks such as shaving (Bien et al., 2003; Chang and Park, 2003), washing the face and hands, brushing teeth, combing hair, shaving, applying makeup (Brose et al., 2010), and dressing (Canal et al., 2021; Nanavati et al., 2024b).
The importance of being able to manipulate a wide variety of objects benefits the ability to get around, for example with an electric wheelchair by being able to move blocking objects out of the way. This is also a task desired by potential users (Nanavati et al., 2024b).
Miscellaneous tasks were mentioned with taking medicine (Nanavati et al., 2024b), working (Nanavati et al., 2024b), leisure activities (Stanger et al., 1994; Nanavati et al., 2024a), turning on and off switches (Bien et al., 2003; Chang and Park, 2003), opening and closing doors (Bien et al., 2003; Brose et al., 2010; Chang and Park, 2003), scratching oneself (Bien et al., 2003; Chang and Park, 2003) and making tea. As sources from 2003 show, changing CDs and removing paper from fax machines are examples of how task preferences change over time (Bien et al., 2003; Chang and Park, 2003). As the accessibility of smart home applications continues to evolve, it is expected that more tasks will become less important for robot assistance, such as flipping light switches or opening doors in the home.
Based on these results, a first selection of tasks was selected for the Wizard of Oz design to present in the studies. In Section 4.1 the literature findings were used to group the responses from the online survey and to compare results related to the change in task prioritization over time in Section 5.1.
2.3. Gaze as control input
The first eye-tracking controllers for robotic arms were developed as early as 2001 (Kim et al., 2001). Since then, several systems have been developed, of which only a few have become commercially available (Fischer-Janzen et al., 2024; Suryadarma et al., 2024; Brose et al., 2010). They differ in the control approach used to guide the robot. A popular design approach is directional control, where the robot moves in the direction of gaze (Alsharif, 2017; Sunny et al., 2021). Graphical User Interfaces (GUIs) are available in which the robot is moved with gaze-selected buttons, taking advantage of more accurate stationary eye-tracking systems (Dragomir et al., 2021; Stalljann et al., 2020).
Going a step further, human intention in eye-tracking control can be predicted from task-related eye movements (anticipatory gaze (Aronson and Admoni, 2020)) based on the natural gaze behavior in hand-eye coordination (González-Díaz et al., 2024). Task anticipation has been explored to estimate task selection from the gaze pattern (Huang and Mutlu, 2016; Wang, 2021). It can be predicted before the action begins and during its execution (González-Díaz et al., 2024; Selvaggio et al., 2021). A crucial step in using gaze as control input is to decompose the behavior into subtasks to obtain specific gaze patterns that can be used to train AI models (Hayhoe and Ballard, 2014). Another application is to use gaze as a complement to generate more dense and accurate information, as is done in Brain-Computer-Interface development (Onose et al., 2012; Ivorra et al., 2018).
Especially reading task intention from gaze allows to use low-level cognitive inputs to realize systems with high autonomy, making it easier to interact with the system (González-Díaz et al., 2024). In contrast to these possibilities, there are some challenges that need to be considered when implementing gaze as a control input:
  • Unintended selection: When gaze is used as a pointing device, e.g. to select an object, rules must be applied to ensure, that not all fixations are selection triggers. This issue is called the Midas Touch problem and describes an unintentional selection of an object (Dünser et al., 2015; Jones et al., 2018; Stalljann et al., 2020). Reasons include unintentional and reflexive eye movements, overlapping objects, distraction, stress, and limited accuracy of the eye-tracking device (Holmqvist and Andersson, 2017).
  • Context misinterpretation: It is caused by overlapping objects or task ambiguity. When objects are close together and the eye-tracker accuracy is not sufficient, the gaze can be tracked on the wrong object leading to faulty interactions between robot and object (Jones et al., 2018). In the interview conducted by Herr et al. (Herr et al., 2016) participants stated: “Well, now I look at you [the interviewer], but I actually focus on the bottle behind you. That would be difficult for the robot to recognize”. This behavior suggests that attention to a particular location is sometimes unrelated to intention or cognitive process.
  • Physiological factors: Physiologically, there are some factors that affect the accuracy of the gaze. Neurological conditions, fatigue, drooping eyelids and strabismus affect the functionality of eye-tracking devices (Holmqvist and Andersson, 2017). The robotic workspace and gaze movements differ in dimensions. Three-dimensional robot workspace has to be covered by two-dimensional eye movements. Eye tracking glasses have decreasing accuracy in the periphery of the field of view. When head movement is limited, as in the case of the stakeholder group, these peripheral regions are more likely to be used to fixate objects, leading to increased instances of contextual misinterpretation.
When such challenges occur, it can lead to frustration and high cognitive load for the user to avert such situations (Dünser et al., 2015; Jones et al., 2018).
2.4. Participation in assistive robot design
Aktion Mensch, a german organization for social projects, states that current assistive devices often fail to meet the needs of users (Borgstedt and Möller-Slawinski, 2023). In eye-tracking controls most studies recruit non-disabled people as participants (Fischer-Janzen et al., 2024). Reasons why people with disabilities are rarely included in studies are recruitment challenges, small sample sizes, and transportation logistics (Nanavati et al., 2024a). As shown by Stalljann et al., control inputs are experienced differently by disabled and non-disabled participants (Stalljann et al., 2020). Including people with disabilities helps in the design process of assistive devices by gaining insight into their lives and needs (Baumeister et al., 2021; Stanger et al., 1994; Nanavati et al., 2024a).
Nanavati et al. (Nanavati et al., 2024a) discusses approaches based on three-dimensional key features to find the most appropriate study design for the goal of the study. The first dimension decides between individual or community-level participation. With individuals, the chance of a deeper insight into the user’s life and preferences can be gained, while community work can lead to a broad insight from many participants. It also affects the time needed to conduct the studies and the connection between participant and experimenter. The second dimension highlights the logistical burden on participants and researchers. People with severe physical disabilities are usually dependent on their caregiver. Depending on the robot setup, it is easier to conduct studies at the participant’s location, to reduce the need of special transportation and availability of caregivers. The third dimension describes the benefits to researchers and the community. Participants have an intrinsic motivation to participate in the design process of assistive devices and robots. They can benefit by providing impetus to adapt the systems to their needs. Researchers benefit by potentially making important connections for future work.
Although it does affect the length of the study, it is worthwhile to involve people with disabilities in the design process. These findings influenced the study design in Section 3.4.
3. Methods
In the introduction and background, we presented influencing factors and user challenges that affect the adoption rates of robotic arm systems. These challenges and factors were partially incorporated into the design of the system, presented in Section 3.1. Several groups of interest were interviewed with two study designs, to obtain broader feedback on the shared control design. The setup described in Section 3.3 and 3.2 was used in both studies. An online survey was sent to stakeholder groups, as well as their families and medical professionals and is presented in Section 3.4. Measurements were conducted to verify technical functionality and the reaction of participants interacting with the robot in person (see Section 3.5).
3.1. Shared control design choices
The initial idea of the system is to include a shared control in which parts of the robot movement are automated, such as trajectory planning, grasping and object manipulation. The level of autonomy is presented following the guidelines of Beer at a. (2014) (Beer et al., 2014).
What task is the robot to perform? This shared control is designed to assist people with severe physical disabilities in everyday tasks. Ongoing work is realizing a framework that can be extended to include everyday activities previously unknown to the robot. The first approach includes the tasks of grasping and handing drinks, filling a glass with a drink, and grasping and placing objects in locations selected by the user. These tasks can be performed in unsorted environments such as at home and in public places, e.g. a restaurant.
These tasks represent two of the most preferred tasks for robotic assistance in everyday life, as shown in Section 2.2.
What aspects of the task should the robot perform? The chosen input modality is eye-tracking, since eye movement is often unrestricted or only slightly restricted in contrast to other motor functions such as arm movement in cases of locked-in syndrome or cerebral palsy. The control is realized by fixating the desired object and the robot decides based on the selected object which available task to perform. Multiple objects can be selected by gaze to make task selection more robust. For example, looking at a cup could indicate that the user wants to drink from it, move it, fill it, or combine available tasks. By adding the ability to combine the cup with other objects, such as selecting a bottle or a location, the task can be distinguished into refilling the cup or placing it in the selected location.
Based on object and ask information, the robot performs the task autonomously. It determines the location of the object, a suitable grasp related to the task, and calculates the optimal trajectory.
To what extent can the robot perform those aspects of the task? The user is always in control of the robot. Emergency stop methods are considered to ensure the well-being of the user. Validation methods are planned to intervene in case of incorrect task execution before the robot starts the task. Settings are planned to reduce the robot’s velocity if desired. With the implementation of dwell time selection of objects, incorrect interaction with the object is reduced with natural gaze behavior, thus minimizing the Midas Touch problem.
What level can the robot’s autonomy be categorized? According to the classification of Beer et al. (Beer et al., 2014) robot autonomy is categorized as Shared Control with Human Initiative.
How might autonomy influence HRI variables? Reliability and functionality of the system must be given. Safety for the user must be ensured and is part of ongoing discussions. This system is designed to be easy to learn and use. Since there is no physical interface such as a display, the correct selection of the object depends on the accuracy of the eye tracker. Objects are selected by focusing on them in real-world settings. If the eye tracker is incorrectly calibrated, the gaze may not be tracked to the object due to high accuracy errors. This raises the question of how serious this error is in this setup, which will be tested with the Wizard of Oz setup. A wearable eye-tracker is used as input device. Influences on social variables such as aesthetics or the occupation of the voice by voice control, which could lead to a rejection of the system, are minimized.
3.2. Wizard of Oz design
Based on the results of the literature review, we decided to do an early evaluation of the design if it meets the expectations for the influencing factors given by the community. This approach has the advantage of shortening the development time by intervening in case of low acceptance or functional challenges. Therefore, the in the following used system represents a Wizard of Oz design, where the autonomously controlled robot tasks are performed by the experimenter, as seen in Figure 1. The objects were placed in a predefined position in the robot’s workspace. All trajectories needed to realize the automated task execution were hard-coded and could be selected by the experimenter clicking on the program. The safety of the participant was ensured by separating the workspace of the participant and robot.
The experimenter was informed about the task sequence of the user and observed the participant’s visualized eye movements in a live video on the monitor. As soon as the participant fixated on an object, the corresponding task was started via the robot’s web interface.
Five tasks were selectable with the Wizard of Oz system: Picking up food with a fork, turning on a light switch, picking up and placing a block, scooping from a bowl, and filling a glass with water and handing it to the participant. Each task was related to different objects, as shown in Figure 2. The selection were based on the initial idea for the system and the outcomes of the literature review.
A video of task performance was created to be implemented in the online survey to visualize the function of the system (Section 3.4). This video is included in the Supplementary Materials. The system was also used in the hands-on study, where participants operated the system in a controlled environment (Section 3.5). Both studies were reviewed and approved by the ethics committee of the first author’s institution.
Figure 1. Wizard of Oz design: The participant selects a task by focusing his or her gaze on the desired object. The gaze is visualized on the computer by displaying the live-video of the eye-tracker. The experimenter starts the task via the robot’s web interface. The robot then performs the task and returns to the home position to wait for further commands. The home position is used as indication that the participant can select the next task.
[…]
Wizard of Oz design: The participant selects an object by focusing on it with his or her gaze. Gaze is visualized on the computer by showing the eye-tracker’s live-video and including a red circle where the user currently looks. Without any further verbal communication the experimenter starts the task via the robot’s web interface by supervising the gaze movement of the participant. Afterwards, the robot performs the task and returns in the start position to wait for further commands. The home position is the indication that the participant can select the next task.
Figure 2. Nine interactable objects were presented with which five tasks were performed. Participants were shielded from visual distractions by green walls surrounding the experimental area.
[…]
The figure shows the setup with nine objects placed either on the table or, in the case of the light switch, on a plywood wall that could be selected to perform five tasks. The scene is seen from the participant’s point of view. The robot was mounted on the opposite side of the table, slightly off-center to the left of the participant. The first task described switching a light switch. The plywood wall with the light switch was on the far left side of the table. The second task, Pick and Place, could be started by selecting either the platform between the light switch and the robot or the block on the far right side of the table. Filling a cup and reaching it to the participant described the third task and was selected by looking at either the cup or the bottle placed next to each other at the same depth of the platform, slightly to the right of the robot. The fourth task was initiated by selecting either the food item placed on a plate in front of and between the cup and bottle, or the fork right next to it. Lastly, scooping food with a spoon was selected by looking at the spoon to the right of the fork or a bowl placed behind the block. Participants were shielded from visual distractions by green walls surrounding the experimental area. The emergency stop button was placed on the table in the front left corner in front of the light switch so that participants could reach it if they felt unsafe.
3.3. Technical setup
The Wizard of Oz setup consists of a robot (Kinova, Gen3 with a gripper Robotiq, 2F-85) controlled via the standard web interface provided by the manufacturer. Data were recorded using wearable eye-tracking glasses (Tobii, Tobii Pro Glasses 3, accuracy = 0.6). Gaze visualization, data analysis, and post-processing were performed using Tobii Pro Lab (Pro, 2014). The control computer to move the robot and visualize the participant’s gaze was placed 2 m behind the participant in the hands-on study so that the user was unaware of the experimenter’s role.
3.4. Online survey
The online survey was chosen as the data collection tool because it provided a standardized way to present information and collect feedback. The purpose is to gain broad feedback from people with disabilities, their family members and caregivers to assess the usability and accessibility of the system. The logistical challenges described by Nanavati et al. (Nanavati et al., 2024a) were minimized by providing an overview of the systems to the stakeholder group via online survey and video. It was advantageous because it allowed immobile people to participate through the use of smart devices. People living at home could participate, as opposed to a real-world demonstration where the authors could only visit a few facilities. In addition, it offered the convenience of completing the survey at the participant’s preferred time, accommodating those with complex schedules or who require assistance. Due to the anticipated limited availability of accessible computer controls, an extension was designed for family members and caregivers to obtain more comprehensive feedback. The purpose of this online survey was to answer the following questions:
  • Are potential users able to use the system?
  • Which tasks are preferred so that the robot has the greatest benefit in everyday life?
  • Were the participants satisfied with the presented system?
3.4.1. Distribution method
The survey was available online from September 18, 2023 to January 22, 2024. 32 organizations in Germany, Switzerland, and Austria were contacted by email to reach the stakeholder groups. With reference to the distribution channels mentioned by Nanavati et al., we selected the following organizations and platforms (Nanavati et al., 2024a). Seven organizations shared the survey, such as FGQ (Paraplegic Support Association), LIS e.V. (Locked-In Syndrome Association), Diakonie (social services), Lebenshilfe (non-profit association), and the Swiss Paraplegic Centre. On social media, the survey was shared a total of three times by the authors’ private accounts and by their institutions. FGQ shared it in their private Facebook group and a post was made in the private forum of the Swiss Paraplegic Centre.
Figure 3. Distribution of questions asked for each group. Q outlines the questions’ context and the number of the questions for each category is given.
[…]
Distribution of questions asked for each group. Q outlines the context of questions asked and the number of the questions for each category is given.
3.4.2. Survey structure
The online survey consisted of 31 questions divided into three groups and a video describing the eye-tracking control of the robotic arm structured as shown in Figure 3. The video showed the calibration process, the light switch task, the pick and place task, and the scoop task. In the survey, people with disabilities were asked questions to assess their level of disability and daily performance. Their need of the control modality, their access to assistive devices, and their evaluation of the presented eye-tracking control was surveyed. Open-ended questions were used to identify tasks that the robot should be able to perform. Family members and caregivers were asked about the amount of time they spend providing care and their estimation of the number of people with physical disabilities in their care facility. In contrast to the questions above, family members and caregivers were first asked what tasks they would like a robot to perform in general, regardless of the type of robot. In this situation, the participants were unaware of the robotic system. Then the system was presented and questions similar to the first questionnaire were asked. Finally, all groups were given the opportunity to provide feedback on the survey and the system itself.
All of the questions asked in the questionnaire, along with the participants’ responses, are available in the Supplementary Materials. The survey was designed with LimeSurvey.
3.4.3. Evaluation method
A frequency analysis was performed to evaluate the users’ perception of the system and the desired tasks. The tasks were categorized as introduced in Section 2.2: tasks related to eating, drinking, picking up objects, getting along, personal hygiene, and miscellaneous tasks.
3.5. Hands-on user study
In addition to the stakeholder evaluation of the system, the Wizard of Oz design was used to analyze real-world interaction. Evidence from the online survey suggests that participants are able to use the system based on their motor function and understanding of the system. The system realization might pose challenges in terms of functionality due to insufficient accuracy or natural gaze behavior. Participants were asked to interact with the robot to record gaze data that would be used to evaluate hardware constraints. Questions answered by this study include:
  • Would the prototype eye-tracking controller work in the controlled environment of the lab?
  • What are the technical difficulties with the design?
  • Does participant perception vary greatly between the hands-on study and the online survey?
3.5.1. Procedure and participants
24 participants took part in the study (7 female/17 male, mean age = 30.5, 11 non-prior experience group/13 prior-experience group with robots) to minimize the influence of learning effects or behavioral adaptation related to the 24 emerging permutations in the task sequence. Each participant was informed and gave consent to be included. They were asked to sit in a chair in front of the robot workspace. The setup and calibration of the eye-tracker was performed as described in (Pro, 2024). Task start was selected by looking at one of two objects in the drinking, eating, and picking and placing tasks (see Figure 1). Participants chose the objects according to their preference. The experimenter performed the role of the automation as described in Section 3.2.
3.5.2. Data analysis
An evaluation of the Areas of Interest (AoI) and the dwell time is necessary to estimate the need for error minimization methods and to determine constraints for the object recognition algorithms. In combination with this technical evaluation of the system, the perception of the non-disabled participants was analyzed. In the Tobii Pro Lab events, Times of Interest (ToI), gaze movements during object selection, and Areas of Interest (AoI) were defined and labeled on the data. Metrics such as dwell times, which describe the total fixation time of an AoI relative to the ToI, and the number of missed hits were calculated by Tobii Pro Lab and exported for further analysis. Based on these parameters, the technical evaluation was performed.
Insights into the factors mentioned in Section 2.3 and Section 2.1 are gained by interpreting gaze data through Times of Interest (ToI), type of gaze movement, and Areas of Interest (AoI) to distinguish dwell times of AoI hits. These parameters can be used to describe the scene based on visual attention. In order to facilitate the understanding of such terms, definitions according to Holmqvist and Nyström (Holmqvist and Andersson, 2017) and Tobii (Pro, 2014) are presented below.
  • Event: Describes in Tobii Pro Lab the time stamp when an event occurs. In this case it marks the start and end time of a task.
  • Time of Interest (ToI): Interval between two related events
  • Area of Interest (AoI): A polygon corresponding to the outline of an object. The resulting area describes a visual stimulus.
  • AoI hit: Describes the entry of the tracked gaze into an AoI.
  • Dwell Time: Describes the time the gaze remains within an AoI after the initial entry. In this study, the gaze was allowed to exit and enter the AoI. It was the sum of time while the gaze stayed within an AoI.
  • Type of eye movement: In this study, fixations and saccades were of most interest. Fixations refer to the time the eyes remain stable and focused on a specific point or area of interest. It was not distinguished from smooth pursuit, in which the head is moved while the gaze rests on a fixed point in the scene. Saccades are rapid changes in gaze direction. They connect fixations. Other eye movements such as vergence movements and microsaccades were not investigated. For further information on gaze movement we refer to Holmqvist and Nyström (Holmqvist and Andersson, 2017).
4. Results
4.1. Online survey
20 out of 71 participants completed the survey. Of the 51 participants who filled out the survey, 36 surveys with no responses (N/A) had to be excluded. If participants completed part of the survey and answered more than page two of seven, the surveys were included in the study. Therefore, 34 complete and partially completed surveys were used in the analysis.
Figure 4. Distribution of responses from online survey participants. Q1 to Q3 were answered by participants with disabilities, Q4 to Q6 by family and friends, and Q7 and Q8 from healthcare professionals.
[…]
This bar chart shows the distribution of responses of online survey respondents. Eight questions are presented that highlight the most interesting findings. People with disabilities were asked the following three questions: First question: Did you have assistance in completing the survey? Responses: fully completed surveys 2 yes, 12 no, 0 not answered. Second question: Do you have access to a robotic arm? Responses: fully completed surveys 1 yes, 13 no, 0 not answered. Third question: Could the system presented be used everyday? Responses: fully completed surveys 9 yes, 2 no, 3 not answered. Family members were asked the following three questions: First question: Do you have access to assistive devices? Responses: fully completed surveys 3 yes, 6 no, 0 not answered. Second question: Are you satisfied with the setup of the system? Responses: fully completed surveys 1 yes, 2 no, 6 not answered. Third question: Would the people you care for be able to use the system? Responses: fully completed surveys 2 yes, 1 no, 6 not answered. Caregivers were asked the following two questions: First question: Are you satisfied with the setup of the system? Responses: fully completed surveys 4 yes, 1 no, 6 not answered. Second question: Would the person you care for be able to use the system? Responses: fully completed surveys 6 yes, 0 no, 5 not answered.
Figure 4 summarizes the results of the survey. 14 participants identified themselves in the group of people with disabilities, eight in family members, and 11 in caregivers. Only once did the respondent not select one of these target groups (”other”). The authors decided to include the participant in the family group because of the affiliation description ”Family but without care share”. The results show that most people with disabilities were able to complete the survey on their own (Q1), leading to the conclusion that they have an assistive device that allows them to use a smart device. This conclusion is based on the participants’ statement that most of them cannot move their hands (10 out of 14).
After viewing the video, participants were asked if they or the people they care for would be able to use the system. More than half of the participants confirmed this (17 out of 32; Q3, Q6, and Q8). Only three participants stated the opposite. The caregivers and family members were also asked if they were satisfied with the setup process (Q5 and Q7). Five stated they were satisfied, while three were not. One reason given was the amount of time it took to set up.
Participants were also asked what tasks they would like to accomplish with the system. The group of people with disabilities gave a total of 14 responses: Eating (3 mentions), drinking (2 mentions), reaching for cutlery, pushing buttons (elevator or electric door), applying makeup or brushing their hair, cleaning, scratching, opening a bottle, and lifting objects of various weights from the floor (2 mentions). Family members and caregivers were asked this question twice. First, they were asked to identify situations in which a robotic system could assist with everyday tasks while they were unaware of the system demonstration. Second, they were asked about the presented system. There were 13 responses to the first question. They rated the following tasks as useful: Making the bed, brushing teeth, assisting with eating, thromboses prevention, putting on socks, documentation, going to the bathroom, doing laundry, cleaning (2 mentions), assisting with walking, assisting with doctor’s appointments and correspondence, silent alarm for service calls, and making phone calls. There were 14 responses to the second question: Eating (5 mentions), drinking, turning on lights, operating the aforementioned lift system, brushing teeth, preparing a drink, opening doors, lifting objects (2 mentions), making phone calls and moving objects. In comparison both questions show a high potential for robotic systems and digitization. In the first case, the focus is on assistance with household tasks and communication, while in the second, user-centered tasks are preferred.
4.2. Hands-on study
4.2.1. Gaze behavior analysis
The analyses of dwell times in relation to object selection events was used to find technical requirements on the eye-tracking controller. Assumptions such as correlations between object size and dwell time duration were investigated for the dwell time trigger anticipated in the design.
Table 1 shows the distribution of participants’ choice of object selection and information about dwell time. A visual attention map is shown in Figure 5. The chosen interval was 1043 ms before the robot started to move and was based on the time steps made by the evaluation software. Both the data in Table 1 and in Figure 5 used the set intervals. In this time period, events such as searching for the object or deciding which object to select were excluded. It was ensured that the participant only focused on the selected object. The average duration of the test was 2 min 41 sec (SD = 9.8 sec).
The average dwell time was calculated to determine the maximum dwell time limit for the controller, including the time from fixation start until the robot moved. The longest possible dwell time for the participant’s gaze in this setting was found to be 5.612 sec. In this interval, the gaze did not leave the AoI. It indicates that dwell time triggers can be set to high durations, but in the following limitations of the maximum dwell time are presented.
In this study, we assumed that the dwell time is longer the larger the object, due to the larger AoI. Accuracy errors would have less weight than in the small AoIs. The correlation between object sizes and corresponding dwell times were measured and are given in Figure 6. The bottle has a higher standard deviation in dwell time than the cup, implying longer dwell times in some participants’ trials. As seen in Table 1 this is also visible in the dwell times above 500 ms. Reasons why the correlation may not apply are, for example, the fixation of the bottle cap, which is closer to the edge of the AoI than the cup handle. The gaze is more likely to leave the AoI, resulting in lower dwell times and indicating the influence of intention. In comparison, the fork and the block have approximately the same average dwell time and SD, despite having different object sizes. Our first hypothesis was extended that the geometry of the object also influences the dwell time. This influence is indicated by the small width of the fork handle, being less than 1 cm. When accuracy errors occur, it is more likely that gaze is tracked outside the AoI.
Table 1. Information about distribution of selection by participants, dwell times and missed AoI hits.
Object Number of selections Occurrences of dwell time over 500 ms Missed AoI hits
Bottle 11 82% 0
Platform 7 87% 1
Cup 13 92% 0
Light switch 24 75% 4
Fork 9 58% 5
Block 14 56% 5
Food item 11 38% 7
Figure 5. Heat map visualizing the attention of participants during task selection (absolute duration), mapped to the snapshot of the experimental setup. Red areas indicate frequent fixations, green less fixations. Areas without coloring indicate no fixations within the observed intervals. Dashed outlines were added manually to better distinguish the edges of the hotspots.
[…]
The figure shows the experimental setup shown in Figure 1. It is overlaid with a heat map visualizing the attention of all participants during task selection (absolute duration). The heat map is visible on the objects and, in some cases, the area around the objects. Task 1: Three hotspots of low and medium intensity are shown on the light switch. There are few fixations outside the light switch towards the center of the experimental setup. Task 2 - Pick and place: The block shows a high intensity hotspot at the top of the block in the left corner toward the center of the experimental setup. The hotspot extends beyond the boundary of the object. A few fixations to the right of the block are visible in the vicinity of the block. The platform shows an elliptical hotspot perpendicular to the height (y-axis) of the figure with low, medium, and high intensity. Few fixations were found outside the platform above the object. Task 3 - Filling a cup: The cup’s high-intensity hotspot is centered on the top part of the handle and extends slightly beyond the object’s outline. The bottle has a vertically elongated hotspot that starts at a third of the height of the bottle from the bottom and continues to the bottle cap, slightly exceeding the objects outlines. Although only low intensities are shown, the highest intensity is at the bottle cap. Task 4 - Picking a food item: The hotspot on the food item is off-center from the food item itself and above the food item. The hotspot is about the same size as the food item and covers approximately half of the food item itself. A second low-intensity hotspot is shown to the right of the food item, between the food item and the fork. The fork has two hotspots at the base of the fork tines and slightly below in 1/6 of the fork handle. Both hotspots extend across the width of the fork.
Figure 6. Correlation between object size, missed hits and dwell time for the selectable objects. Dwell time is given as the total duration of gaze within the objects AoI 1043 ms before the robot started moving. The line in the box plots marks the median, confidence intervals are given at 95%.
[…]
This graph shows the correlation between object size, missed hits and dwell time for the selectable objects. Dwell time is given as the total duration of gaze within the objects AoI 1043 ms before the robot started moving and visualized as a box plot. In the following the average, median whiskers and outliers are described for each object. Bottle: Median = 983 ms, SD = 826 ms to 1043 ms, 2 outliers at 375 ms and 24 ms. Cup: Median = 1027 ms, SD = 917 ms to 1040 ms, 2 outliers at 715 ms and 113 ms. Fork: Median = 654 ms, SD = 0 ms to 1043 ms, no outliers. Food Item: Median = 0 ms, SD = 0 ms to 1040 ms, no outliers. Platform: Median = 1031 ms, SD = 0 ms to 1043 ms, no outliers. Block: Median = 684 ms, SD = 0 ms to 1043 ms, no outliers. Light switch: Median = 1016 ms, SD = 0 ms to 1043 ms, no outliers. Two line plots representing object size and missed hits were superimposed on the box plot. Line plot results for the object size in square meter: Bottle: 0.018, cup: 0.00657, fork: 0.00475, food item: 0.001225, platform: 0.008649, block: 0.0016, light switch: 0.0064. Line plot results for missed hits of the AoI of each object: Bottle: 0, cup: 0, fork: 5, food item: 7, platform: 1, block: 5, light switch: 4.
The influence of gaze direction was found by examining the low dwell times of the food item placed closest to the user. Looking at the gaze distribution in relation to the scene camera coordinate system, as shown in Figure 7, it is evident that participants tended to move their head to center their gaze when looking at a peripheral object to the right or left. In contrast, participants tended not to move their head when looking down. The detected gaze location was less accurate, due to higher scatter closer to the eye-tracker’s periphery. Physiological impact cannot be excluded such as shadowed pupils by the eyelashes. This resulted in seven missed AoI hits for the food item.
Significant correlation was investigated between dwell time, object size, and gaze direction towards the object. The Shapiro-Wilk test (H = 0.771, df = 97, p = ¡0.001) and the Levene test (H = 2.658, df1 = 6, df2 = 90, p = 0.02) showed no normal distribution or homogeneity. A Kruskal-Wallis test was applied and showed significant differences ($H(6)=17.442$, $p<0.008$). Post-hoc analysis using Dunn’s test with Bonferroni correction revealed significant differences ($p<0.05$) between the following pairs: food item and bottle, food item and platform, food item and light switch, food item and cup, and block and light switch. These results indicate that the measured data for the food item are significantly different from the other objects. Spearman’s $\rho$ was used to analyze the correlation between object size, object position, and dwell time. The frame in Figure 7 was evenly divided into five-by-five sectors in horizontal and vertical directions, so that gaze location could be described by a coordinate map. The analysis showed that dwell time and object size ($\rho=0.373$, $p<0.001$) as well as dwell time and offset in y-direction ($\rho=0.263$, $p=0.009$) were positively correlated. The correlation between the offset values in y-direction and x-direction is weaker and does not correlate with dwell time ($\rho=0.160$, $p=0.116$). This indicates that the assumption that the dwell time is influenced by gaze location towards the object and the object size is correct.
Figure 7. Left: Exemplary representation of the viewing directions to explain the resulting heat map. Visualization of the viewing point in relation to the visual field of the scene camera. Right: Heat map of gaze position during task selection. The field of view is divided into 5 by 5 fields indicating a 5-section coordinate map used to determine the correlation between dwell time, gaze location, and missed AoI hits. Gaze direction was mostly centered and downwards while task selection.
[…]
This figure shows the heat map visualization of gaze position in the field of view of the eye-tracking camera during task selection. The field of view is divided into 5 by 5 fields indicating a 5-section coordinate map used to determine the correlation between dwell time, gaze location, and missed AoI hits. Starting from the upper left corner, the fields are labeled as follows: A1 is the upper left corner, A5 is the upper right corner, E1 is the lower left corner, and E5 is the lower right corner. The highest intensity of the participants’ attention is found in C3, D3, and E3. Medium intensity is also found in C2, D2, E2, D4, and D5. Low intensity is found in all fields except row A and columns 1 and 5. Therefore, the participants’ gaze is centered in the middle of the field of view in the x-direction and mostly in the middle and lower part in the y-direction.
4.2.2. Perception of hands-on application
In general, satisfaction with the system was expressed in the hands-on study. Participants responded with following comments: 9 positive, 3 negative, 9 times positive and negative, and 3 times with no comments. All statements can be found in Appendix A. Differences were found between the responses of people who work regularly with robotic arms (called prior-experience group below) and people who have no experience with robotic arms (called non-prior-experience group). The non-prior-experience group gave mostly positive comments and focused more on the description of tasks, while the prior-experience group focused on robot kinematics, robot behavior, and user safety, which partially included negative aspects.
While the non-expert group reported being interested, entertained, satisfied, and impressed, they also perceived confusion and uncertainty. The task design itself caused confusion, as participants were not sure how to react to the glass presented by the robot. Considering a real-life application of filling the glass, one of the participants expressed fear of overfilling the glass. Regarding the robotic system, they were satisfied with good reaction time and intuitive control. The velocity was ambivalent, with some participants saying it was just right and others saying it was too slow. Importantly, users reported feeling safer because the robot’s movements were limited to the edge of the table. The fork movement may have had an unsettling effect on the person, as they reported that the fork was not close enough to eat or at the right height for their mouth.
The prior-experience group comments were mostly negative about the robot’s trajectory and the measurement setup. For example, the gripper could crush someone’s hand if the user reached into the setup, due to the lack of pressure sensors. The placement of the emergency stop was a concern for the well-being of future users due to its distance from the participant. The shaky movements of the robot, the non-optimized trajectories, and in one case the robot accidentally touching the bottle while moving to the next object were mentioned. The perceived motor vibration was uncomfortable and the sound of the dropped fork was too loud. Reaction time and object recognition were mentioned as positive. In addition, two participants from both groups stated that they felt uncertain that the robot would stop the task if they looked away or at other objects while the robot was performing the task.
5. Discussion
In Section 5.1 two of three usability questions and in Section 5.2 two of the three functionality questions are answered asked in the Methods. In Section 5.3 the last question of both studies will be discussed, which combines the perception of the online study and hands-on study. The resulting adaption features to improve the initial design are given in Section 5.4.
5.1. Usability
5.1.1. Are potential users able to use the system?
Yes, the survey showed in Q3, Q6, and Q8 (Figure 4) that most participants are able to use the system based on the video presentation (Answers: 17 yes, 3 no, 14 N/A). The high amount of N/A results were a result of including partially filled out surveys.
In the open-ended questions, statements were made that the participants would like to have more information about the system to better evaluate the maintenance needed and the function of the system. These requests did not imply that the participant would change his or her mind after receiving the additional information. One participant stated “I wonder if the system can be used by people with intellectual disabilities due to their reduced cognition”. This underlines the importance of considering a variety of disabilities in the design process of such systems in order to make them accessible to a wide range of users. It highlights again the necessity to involve the community in the design process.
5.1.2. Which tasks are preferred so that the robot has the greatest benefit in everyday life?
The categories of eating, drinking, environmental interactions, pick and place, personal hygiene, and miscellaneous tasks were found in the literature and are listed in Section 2.2. The participants’ responses from Section 4.1 were sorted into these categories that lead the distribution in Figure 8. The most requested task in the online survey that the robot should perform is assisting to eat or drink, followed by environmental interaction, which includes turning lights on and off, pressing buttons, opening doors, and pick and place objects. Picking up and moving objects was mentioned in different contexts such as moving heavy objects out of the way. Personal hygiene consisted of a variety of different tasks including doing makeup and hair, brushing teeth, and cleaning. Miscellaneous tasks included making phone calls and scratching themselves. These results can be corroborated with other findings (Chang and Park, 2003; Stanger et al., 1994; WHO, 2010). The variety of tasks expressed shows that participants were not biased by the tasks shown in the video.
Figure 8. Classification of all mentioned tasks in relation to the presented system. 28 tasks were classified into six categories containing the tasks mentioned by the participants.
[…]
This pie chart shows the classification of all mentioned tasks in relation to the presented system. 28 tasks were grouped into six categories containing the tasks mentioned by the participants and are shown in percentage. The largest category is nutrition with 52%, the second largest category is everyday tasks with 19%, the third is mobility with 11%, the fourth is two categories with 7% each described by the terms personal hygiene and well-being. Finally, tasks classified as communication are mentioned with 4%.
Brushing one’s hair, applying makeup, or scratching are everyday tasks that require close human-robot interaction while ensuring the safety of the user. The shared control approach presented would require face and body recognition if the user is unable to move their head to a certain position. Ethical questions arise when personal data is recorded, including physical and physiological appearance. Current discussions in this area suggest that the user must be informed about the system functions and the personal data collected. This would allow the user to decide whether the privacy limitations or the benefits of the system outweigh (Fosch-Villaronga et al., 2020). The aim must be to present this information in simple language so that people with physical and cognitive disabilities can understand it and make decisions.
5.2. Functionality
5.2.1. Would the prototype eye-tracking controller work in the controlled environment of the lab?
Initial testing showed that 70% of object selections would result in selection success when the dwell time trigger is set to 500 ms. Although, dwell times can be shortened to enhance selection success as shown in other literature (Holmqvist and Andersson, 2017), this could lead to an unintentional object selection as described in Section 2.3. An appropriate threshold must be explored with the prototype system.
Accuracy issues are one of the main reasons why gaze was tracked outside of an AoI. Therefore, influences on AoI misses and dwell time were investigated in Section 4.2. Four dependencies were found:
  • Object size: As shown in Figure 6, object size correlates significantly with dwell time. Small objects like keys are part of everyday activities. Excluding such objects from interaction, would lead to major drawbacks in participation in daily life. The available workspace is limited by the length of the robotic arm. This also limits the relative size of the object in the camera scene. With this limit, the resulting accuracy error can be calculated and checked with the desired interactable object size to estimate if this would cause problems. In this case, the accuracy of the eye-tracker is $\Delta_{deg}=0.6$°. With an effective workspace of $d=89.18$ cm for the used robot arm (inc., 2022), the maximum accuracy error is 9.34 mm ($\Delta=d*tan(\Delta_{deg})$). Since all selected objects were larger than 9.34 mm, it is assumed that the prototype would work.
  • Geometry: The AoI of the fork could lead in a failed selection due to its width. Currently, there are two approaches in planning for estimating object outlines. Object recognition can be used to create bounding boxes, solving the issue due to the new geometry. Overlapping of bounding boxes may occur more often, leading to unintentional object selection and context misinterpretation. Object segmentation can lead to boundaries which fit to the object’s outlines. In this case the geometry of the object would be traced with high accuracy, solving the bounding box issue, leading back to the initial width issue.
  • Gaze location towards the object: In the eating task example, participants tend to move their head to the left and right instead of up and down (see Figure 7). The resulting higher accuracy error was due to the glasses periphery. Considering physical disabilities of potential users, an immobility of the head cannot be excluded.
  • User intention: Participants were instructed to focus their gaze on the selected object, but not on which region in detail. In the case of the bottle, the task was to fill the glass. The bottle needs to be opened first, which explains why the bottle cap was fixated, indicating the hand-eye coordination and intention behind the task as discussed in Section 2.1.
5.2.2. What are the technical difficulties with the design?
Ambiguous object selection can occur due to object overlap and was found in the case of the fork (see Figure 5). In addition to the accuracy challenges, robust object selection may be the most difficult challenge in implementing the prototype.
Task validation methods and emergency stops must be ensured, as found in Section 2.1. Approaches such as looking into the corner of the field of view could be challenging to interpret as an emergency stop because it is processed as the same data type as prolonged blinking in the eye-tracking software. Another approach describes the inclusion of gaze gestures (unique sequence of eye movements (Alsharif, 2017)). It needs to be verified whether such a cue would be able to perform for a person in an emergency situation. For task validation, the system needs to be enhanced with feedback methods to communicate the expected task by the robot.
5.3. Differences between Online Study and Hands-On User Study
5.3.1. Were the participants satisfied with the presented system?
In both studies, participants were mostly satisfied with the system. The challenges that were uncovered are summarized in Section 5.4.
5.3.2. Does participant perception vary greatly between the hands-on study and the online survey?
The unstructured in-person interview with participants in the hands-on study and the open-ended questions in the online survey, provided more diverse feedback on this shared control design. The results are consistent with and extend current literature in this area. It adds information for the design of gaze-controlled assistive robotic arms. Perceptions of the system do not vary but, highlight different aspects of the design.
Involving the stakeholder group yielded insights about everyday tasks and their daily life with their disability. The prior-experience group revealed potential drawbacks in the technical design process, such as concerns about safety, kinematic selection, and trajectory planning. Although the non-prior-experience group did not have in-depth feedback on certain aspects of the life with a disability or technology, they helped build a larger pool of participants to measure gaze behavior. In addition, their perception of the robot was more curious and at the same time hesitant than that of the prior-experience group, due to the novelty of robot interaction. Since robotic arms are rarely available to people with disabilities, it is reasonable to assume that in a real-world test, people with disabilities would respond in a initial robot interaction similarly to the non-prior-experience group.
5.4. Adaptation of the initial Shared Control design
Based on these results, the design can be refined to meet the participants’ expectations of the system. The next step is to automate the experimenter’s control task so that the robot can autonomously perform tasks selected by the user. The framework must have the following features to realize the system:
  • Gaze detection on objects: Due to the outcomes in Section 5.2, several approaches are followed to detect objects in the scene camera of the eye-tracking glasses. Unintentional selection and context misinterpretation will be estimated with the prototype.
  • Task availability: The survey showed that the desired tasks were in line with the participants’ expectations. Eating was the most frequently mentioned task. However, handling a fork is a safety risk, when the robot manipulates the fork autonomously and has to be handled appropriately.
  • Task selection: The expected prototype will implement objects associated with multiple tasks. Testing has shown that if the dwell time is set correctly and the algorithm is designed to wait for multiple objects to be selected, the task should be robustly selectable. Dwell times of 500 ms and less were found to be appropriate. As mentioned in Section 2.1, feedback strategies need to be evaluated and integrated into the control to ensure the user.
  • Trajectory, grasping and collision avoidance: The robot has to perform the task by itself. In the current version, the robot’s trajectory was programmed and the objects were placed at the same location for each test. In a real-world scenario this would not meet the user’s expectations. Solutions were found as presented in Section 2.3) and depend on the approach chosen for object gaze detection.
  • Adaptation: On behalf of the ambiguous results of the robot’s velocity perception, an adaptation setting will be included in the prototype so that users can adapt the system to their liking such as robot velocity.
6. Limitations
Additional questions: The safety concerns were not reflected in the responses to the online survey. Including questions such as ”Would you feel safe interacting with the robot?” would provide a more comprehensive understanding of this topic. However, other studies show that potential users are more likely to accept the risks of using the system than to reject it and become dependent on caregivers or family members (Baumeister et al., 2021). This can also be seen here, as the robot-experienced group is aware of the risks of human-robot interaction but is not fearful of interacting with the robot. Specific Wizard of Oz design: The Wizard of Oz system is an eye-tracking-driven design in which objects are selected without the need for a GUI. This approach is uncommon since prototype systems face challenges such as interpreting the desired task, error estimation due to head movements, and a lack of feedback modalities (Fischer-Janzen et al., 2024). Researchers using these adaptation suggestions should check whether the findings can be applied to different input designs. Generalized approach: This approach was not tailored to specific disabilities, so it may lack details necessary to adapt to individual users’ needs. Future work can use this generalization to more easily locate these demands.
7. Conclusion
The goal of this work was to obtain feedback on an eye-tracking-driven shared control for Assistive Robotic Arms to identify community expectations, desired tasks, and technical challenges. The factors for a successful implementation of a shared control were identified and presented. A Wizard of Oz design was created prior to prototyping, as suggested by other researchers.
Based on this design, two studies were conducted. An online survey designed for people with disabilities, their families and friends, and caregivers provided feedback on the community’s expectations for the system, as well as the desired task the robot should be able to perform. A hands-on study with participants with and without robotics experience provided insight into the technical requirements and feedback differences between the video presentation available for the online survey and the real-world application. Gaze data was recorded to measure dwell times and influence dependencies related to the eye-tracking control.
Dependencies for the design of eye-tracking-driven robot controls were uncovered, ranging from user demands on task prioritization, to accuracy limiting factors in eye-tracking. As a consequence, improvements to the initial idea were presented to facilitate the usage of the system.
Acknowledgements.
Both studies were approved by the ethics commission of the Offenburg university of Applied Sciences. The study designs were checked to comply to the ACM Publications Policy on Research Involving Human Participants and Subjects. The Tobii Pro Glasses 3 was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number INST898/33-1. Further, we would like to thank Atikkhan Faridkhan Nilgar and Karl Tschurtschenthaler for the great discussions in statistics and eye-tracking evaluation related to this topic. We would like to thank Katrin-Misel Ponomarjova to proof read the translations of the questionnaire. Lastly, we want to thank Kim Zähringer for her assistance with the robot and video making.
Appendix A Participant statements of hands-on study
In the following the statements of participants of the hands-on study are listed. All statements were translated by the authors to english. Participants without any statement to the system were not listed below.
Table 2. Statements of participants. Prior-experience group: Exp., Non-Prior-experience group: Non-Exp.
P Exp./ Non-Exp. Statement
1 Non-Exp. By noting that the robot does not go over the table, P1 was reassured. “apprehension, causious” system if not had stated, Glass’s task: fear of overspilling and dropping the cup in the real task.
3 Non-Exp. Food task: Lifting a fork so high that you only had to grab was cool. Light switch was hit centered, that was cool. The handle of the cup was held towards the person. That was very practical if you wanted to grab it.
4 Non-Exp. Fork task: Test person was confused because the fork pointed to a fixed point and did not go in the direction of the mouth. General: Confused as to whether the robot cancels the task when looking at another object.
5 Non-Exp. Movements too slow for everyday use. However, it was satisfactory for an initial test.
6 Non-Exp. Food task: Serving food was interesting. Glass task: The glass slipped a little when I put it down. There was no straw in the glass, so it wasn’t clear why the robot handed the glass flat.
7 Non-Exp. Open gripper is uncomfortable when it points at you. Velocity was good.
10 Non-Exp. General: By saying that the robot does not reach over the table, the action was not scary.
11 Non-Exp. General: Responds quickly, which is good. Intuitive operation. Once you know where the robot stops it is less intimidating/scary.
13 Exp. Fork task: First close the gripper a little before gripping the fork completely, less risk of crushing the hand. General: Emergency stop is poorly positioned.
14 Exp. Kinematics moves spongily, drinking task: Kinematics has touched the bottle.
15 Exp. When the fork is dropped, a loud unpleasant noise is heard from the drop.
16 Exp. P16 asked whether tasks are aborted if you accidentally look at other objects. Waiting time is appropriate until the robot moves. It does what you want (intuitive control). Kinematics vibrate during movement, which is negative.
17 Exp. Robot is shaky, jerks. Selection process: Fear of triggering the wrong action when looking at the wrong object.
18 Exp. Movement was a little too fast to the light switch
22 Exp. With pick and place, it was assumed that either the block or the platform could be gripped. Accordingly, P22 was confused.
23 Exp. Robot trembles, trajectory not optimal, bottle badly twisted while handled.
24 Exp. Fork was recognized well, even though you looked briefly at the bottle during the time you had to aim. The pause between aiming and starting the pick and place task was too long!

### v2

@@
 Abstract.
-Advances in eye-tracking control for assistive robotic arms provide intuitive interaction opportunities for people with physical disabilities.
-Shared control has gained interest in recent years by improving user satisfaction through partial automation of robot control.
-We present an eye-tracking-guided shared control design based on insights from state-of-the-art literature.
-A Wizard of Oz setup was used in which automation was simulated by an experimenter to evaluate the concept without requiring full implementation.
-This approach allowed for rapid exploration of user needs and expectations to inform future iterations.
-Two studies were conducted to assess user experience, identify design challenges, and find improvements to ensure usability and accessibility.
-The first study involved people with disabilities by providing a survey, and the second study used the Wizard of Oz design in person to gain technical insights, leading to a comprehensive picture of findings.
+Eye-tracking-driven controls for assistive robotic arms provide people with severe physical disabilities with intuitive interaction opportunities.In this context, shared control can improve acceptance of the robot by partially automating task execution.
+Based on recent literature, we present a Wizard of Oz design for shared control driven by eye-tracking.
+This approach allows for the rapid exploration of user expectations to inform future design iterations.
+Two studies were conducted to assess user experience, identify design challenges, and determine ways to improve usability and accessibility.
+The first study involved people with severe motor disabilities (PSMD) and consisted of an online survey.
+The second study aimed to gain technical insights through direct robot interaction to provide a comprehensive overview of the findings.
 Assistive Robotics, Eye-Tracking, Robot Manipulation, Assistive Device, User Experience, Disability, Wizard of Oz
-ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Pointing devicesccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Development frameworks and environments
+copyright: rightsretainedjournal: THRIjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3796524journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Pointing devicesccs: Human-centered computing Empirical studies in accessibilityccs: Software and its engineering Development frameworks and environments
 1. Introduction
-People with physical disabilities such as locked-in syndrome, cerebral palsy, late-stage amyotrophic lateral sclerosis, or multiple sclerosis benefit from non-haptic controls such as Brain-Computer Interfaces and eye-tracking controllers.
-Implementation as control inputs for assistive robotic arms can enhance user independence, self-determination, and privacy in daily life by reducing the dependency on family members and caregivers (Fischer-Janzen et al., 2024; Awuah et al., 2024; Baumeister et al., 2021).
-Regardless of the controlled technology, challenges have been identified with such input devices in the areas of safety, availability of participants for clinical trials, privacy, interdisciplinarity, and cost (Maiseli et al., 2023; Suryadarma et al., 2024; Fischer-Janzen et al., 2024; Nanavati et al., 2024b).
-This has led to the development of a variety of robot control methods that differ in terms of input modalities and control strategies (Mohebbi, 2020; Fischer-Janzen et al., 2024).
-Depending on the input modality, the achievable varying level of robot autonomy yields different advantages and disadvantages.
-For instance, direct control of the robot leads to the need for mode switches to cover the 3-dimensional space the robot can move, or to control all the joints of the robot, enhancing the needed concentration on the tasks (Fischer-Janzen et al., 2024; Kronhardt et al., 2023; Herlant et al., 2016).
-The implementation of a shared control approach, in which tasks are partially automated, has been demonstrated to effectively address this issue.
-Such approaches have been shown to positively impact interaction perception and reduce the time required to complete tasks (Bhattacharjee et al., 2020; Kronhardt et al., 2023).
-A potential challenge associated with a high level of automation is user discomfort, which can arise if users feel that they are losing control over the system.
-In this context, community feedback is essential to improve design and evaluate usability and accessibility (Stalljann et al., 2020; Baumeister et al., 2021; Nanavati et al., 2024a).
-It provides firsthand insights into user expectations and perceptions of their daily lives, thus enhancing acceptance rates (Nanavati et al., 2024a; Canal et al., 2021).
-This work incorporates feedback from the community on an initial shared control design to identify challenges and improve the design.
-A Wizard of Oz design of the eye-tracking-driven control was developed, with which two studies were conducted.
-The first study was an online survey distributed to people with disabilities, their family members and friends, and caregivers.
-The objective was to assess the usability in daily life and to identify the tasks that participants would prefer to perform.
-The stakeholder group was defined as individuals with physical disabilities who are unable to use their extremities to control the robot.
-In the second study, non-disabled participants with varying levels of robotics experience interacted with the system to obtain technical parameters that would provide constraints for a future control algorithm.
-In this context technical parameters describe for example the maximum and minimum dwell time and the level of accuracy needed for a robust control.
+People with severe motor disabilities (PSMD) often rely on hands-free control modalities, such as Brain-Computer Interfaces, sip-and-puff systems, and eye-tracking interfaces.
+These technologies have been successfully used by individuals with conditions such as locked-in syndrome, cerebral palsy, late-stage amyotrophic lateral sclerosis, or multiple sclerosis.
+As control inputs for assistive robotic arms, they can support independence, self-determination, and privacy in daily life, reducing reliance on family members and caregivers (Fischer-Janzen et al., 2024; Awuah et al., 2024; Baumeister et al., 2021).
+A shared control allows users to give the robot high-level commands, such as selecting an object or goal, by automating task sequences.
+In eye-tracking-driven approaches, advances in AI allow us to move beyond manual mappings toward intelligent interfaces that interpret gaze patterns to extract task intent (Aronson et al., 2018; Belardinelli, 2024; González-Díaz et al., 2024).
+Environmental context, such as the presence of objects, can further improve autonomous decision-making (Zhang et al., 2025; Yang et al., 2021).
+These advances have led to the development of various robot control methods for Human-Robot Interaction (HRI) that differ in terms of control strategies (Mohebbi, 2020; Fischer-Janzen et al., 2024).
+With a higher level of robot autonomy, challenges such as controlling three-dimensional robot movement with two-dimensional eye movements can be resolved.
+Bypassing the need for mode switches, as required for direct control, to cover the 3-dimensional robot space, or independent joint control, can reduce workload and decrease the concentration required for the task (Fischer-Janzen et al., 2024; Kronhardt et al., 2023; Herlant et al., 2016).
+Shared control approaches have been shown to positively impact interaction perception and reduce task completion time (Bhattacharjee et al., 2020; Kronhardt et al., 2023).
+A challenge associated with a high level of automation is maintaining the sense of agency.
+Community feedback is essential to improve design and evaluate usability and accessibility (Stalljann et al., 2020; Baumeister et al., 2021; Nanavati et al., 2024a), as well as to uncover influences and design controls with an appropriate level of autonomy.
+It provides firsthand insights into user expectations, thus enhancing acceptance rates (Nanavati et al., 2024a; Canal et al., 2021).
+The novelty of this work is characterized by a detailed look at usability and functionality regarding robot acceptance parameters in respect to articulated robots used to assist in daily life.
+Current task intent recognition methods are based on object labeling done by object detection or segmentation, which we evaluated in this context on feasibility.
+This work incorporates community feedback on an initial shared control design to identify challenges and guide improvements.
+A Wizard of Oz design of the eye-tracking-driven control was developed and used in two studies.
+The first study was an online survey distributed to PSMD, their families, friends, and medical professionals.
+The goal of the study was to assess everyday usability and identify task preferences.
+The stakeholder group was defined as PSMD who experience significant motor impairments in their arms and torso.
+Previous research has shown differences in the perception of virtual versus real robot interactions (Petrich et al., 2022).
+A second study was conducted, in which participants without motor impairments and with varying levels of robotics experience interacted with the physical system.
+This study focused on deriving technical characteristics that can inform the constraints of a future control algorithm, such as feasible dwell times and accuracy requirements.
 The main contributions of this work are as follows:
   1. (1)
-    Presentation of a shared control design for everyday tasks.
+    Presentation of an initial eye-tracking driven shared control design leveraging object detection algorithms for daily tasks.
   2. (2)
-    An evaluation of functionality and usability through opinions of potential users, their families, caregivers and non-disabled participants.
+    A functionality and usability evaluation, incorporating feedback from potential users, their families, medical professionals, and non-disabled participants with and without prior robot experience.
   3. (3)
-    Improvements in the initial shared control design gained from the studies.
+    Refinements and task preferences for the initial shared control design, derived from insights obtained in both studies.
 2. Background
-2.1. Shared control for robotic arms
+2.1. Gaze as control input in shared controls
 ”Shared control in assistive robotics blends human autonomy with computer assistance, thus simplifying complex tasks for individuals with physical impairments” (Goldau et al., 2024).
-Advances in sensing, inference, modeling, and learning methods have lead to novel developments in the research area (Selvaggio et al., 2021).
-In this context, researchers are investigating impact factors on shared control.
-Literature has been reviewed to find such factors influencing human-robot-interaction.
-three categories were introduced to facilitate the overview over the found outcomes: Field of use, technical challenges, level of autonomy, and user individuality and preferences.
-The field of use establishes the boundaries within which shared controls are to be utilized.
-Robots can be used for daily tasks in unstructured environments (Stanger et al., 1994), household chores (Herr et al., 2016), and personal tasks such as personal hygiene or medical procedures (Wang et al., 2020).
-This determines the positioning of the robot, whether it must be attached to an electric wheelchair, on the bedside, or on a table (Stanger et al., 1994), and leads to the initial setup configuration and limitation in robot choice.
-The technical challenges include reliability and functionality of the system: Safety, robustness and success rate in e.g.
-task completion is a standard that has to be ensured (Stanger et al., 1994; Wang et al., 2020) and the requirements are defined in ISO 13482 (für Normung e.V., DIN).
-The system must be robust to changing or unorganized environments, especially when the robot is mounted on an electric wheelchair (Selvaggio et al., 2021; Stanger et al., 1994).
-The system should be easy to learn and use, which also depends on the complexity of the task (Stanger et al., 1994; Shafti et al., 2019; Goldau et al., 2024).
-Visualization with an interface can help (Shafti et al., 2019; Nanavati et al., 2024b).
-If interfaces in the form of displays are not available other feedback methods can ensure the ease of use by communicating information to the user and reducing uncertainty in using the robot (Shafti et al., 2019; Nanavati et al., 2024b).
-Autonomous robotic task learning, for example imitation learning, ensures the availability of a large range of tasks (Stanger et al., 1994; Selvaggio et al., 2021).
-The availability of general tasks adapted to the field of usage eases the distress in the lack of capabilities (Brose et al., 2010).
-The appearance and aesthetics can influence the results of user’s system perception.
-In the past, connections were found that due to insufficient aesthetics, users expected reduced social communication and isolation (Stanger et al., 1994; Brose et al., 2010).
-Ongoing debates revolve around the level of autonomy and can vary from teleoperated systems to fully autonomous robots (Beer et al., 2014).
-Recent studies have shown that more autonomy is not always better for the user (Nanavati et al., 2024b; Bhattacharjee et al., 2020).
-Autonomous behavior of the robot ”allows reducing the human operator’s workload when performing a task that can be repetitive and/or requires skills, effort or precision levels that exceed those of a human” (Selvaggio et al., 2021).
-If the robot produces more errors in task execution with a high level of autonomy, users tend to reject the system (Bhattacharjee et al., 2020).
-It can further lead to lost trust or a sense of lost control with unpredictable behavior of the system (Nanavati et al., 2024b).
-Validation functions communicating and verifying the next steps with the user can ease this issue (Bhattacharjee et al., 2020).
-On the other hand, the user’s attitude towards a task influences the wish for a higher level of autonomy or a complete delegation of the task to the robot, e.g.
-if the user chooses not to perform the task (Herr et al., 2016).
-In eye-tracking controls these levels of autonomy can be distinguished in teleoperated system, directionally controlled systems and systems with automated trajectory or grasp planning.
-This variation strongly influences the design process (Fischer-Janzen et al., 2024).
-Directional gaze control is easy to learn, provides a sense of control, and reduces cognitive load (Dragomir et al., 2021; Dziemian et al., 2016).
-On the other hand, it is more time consuming to control each of the robot movements separately.
-Shared controls are mostly controlled by the user selecting the object via fixation, indicating the wish to interact with (Catalán et al., 2017; Zeng et al., 2017; Tostado et al., 2016).
-The user’s individuality and preferences vary with personal metrics such as the user’s age, level of disability, living arrangements, and familiarity with technology and affect the system’s performance perception (Stanger et al., 1994).
-The adaptation of the system to user preferences is mentioned by various authors (Bhattacharjee et al., 2020; Selvaggio et al., 2021).
-The ability to adapt the system gave rise to Shared Autonomy approaches, ”where the robot is capable of seamless adaptation of its autonomy level based on its own understanding of the human actions/intentions and of the surrounding environment” (Selvaggio et al., 2021).
-Incorporating the preferences of the user influences the perception and user satisfaction of the system (Cakmak et al., 2011; Canal et al., 2021).
-An example is the participant’s preference for handing an object with a particular orientation (Cakmak et al., 2011).
-A common factor, as known from the use of PCs, is the adaptation of the system to the user’s preferences by offering a settings menu.
-Parameters such as the velocity of the robot should be adaptable (Bhattacharjee et al., 2020).
-Another factor is that offering a variation of input modalities (e.g.
-head or eye-tracking control instead of haptic control) makes it possible to tailor the system to the user’s needs (Bhattacharjee et al., 2020; Nanavati et al., 2024b).
-Users want their interactions with the robot ”to be customized to their impairment, their preferences, and their context” (Nanavati et al., 2024b).
-Other impact factors are a low maintenance system to ensure a reduced workload for others, such as family or caregivers (Bhattacharjee et al., 2020).
-Related to market introduction are factors such as the initial cost and the maintenance cost of the system (Stanger et al., 1994; Brose et al., 2010).
-These may have less impact on the outcomes of the user survey, but are an important factor in the design of the system.
-Due to the closeness between human and robot these technical and ethical challenges coexist (Canal et al., 2021).
-In designing a well-balanced shared control, a trade-off must be found for all factors relating to technical and ethical challenges.
-In the following, we will present a solution that considers and balances the different influences.
-2.2. Task preferences in assistive robot manipulation
-Among the many daily tasks, research has identified the areas where people with physical disabilities prefer assistance.
-This can have significant impact on the design of the system, through the adaptation of grippers or, in the case of shared control, task automation.
-The four most frequently mentioned tasks that an assistive robot should perform, were food related tasks (serving a meal, eating) (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Canal et al., 2021; Bhattacharjee et al., 2020; Nanavati et al., 2024a), drinking related tasks (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003), picking up objects (Bien et al., 2003; Stanger et al., 1994; Brose et al., 2010; Chang and Park, 2003; Nanavati et al., 2024b), and personal hygiene (Stanger et al., 1994; Brose et al., 2010; Nanavati et al., 2024b).
-Picking up an object was further related to the location or a manipulation that was desired with the object such as picking it up from the floor or from a shelf, bringing it near, or moving it.
-Personal hygiene included various tasks such as shaving (Bien et al., 2003; Chang and Park, 2003), washing the face and hands, brushing teeth, combing hair, shaving, applying makeup (Brose et al., 2010), and dressing (Canal et al., 2021; Nanavati et al., 2024b).
-The importance of being able to manipulate a wide variety of objects benefits the ability to get around, for example with an electric wheelchair by being able to move blocking objects out of the way.
-This is also a task desired by potential users (Nanavati et al., 2024b).
-Miscellaneous tasks were mentioned with taking medicine (Nanavati et al., 2024b), working (Nanavati et al., 2024b), leisure activities (Stanger et al., 1994; Nanavati et al., 2024a), turning on and off switches (Bien et al., 2003; Chang and Park, 2003), opening and closing doors (Bien et al., 2003; Brose et al., 2010; Chang and Park, 2003), scratching oneself (Bien et al., 2003; Chang and Park, 2003) and making tea.
-As sources from 2003 show, changing CDs and removing paper from fax machines are examples of how task preferences change over time (Bien et al., 2003; Chang and Park, 2003).
-As the accessibility of smart home applications continues to evolve, it is expected that more tasks will become less important for robot assistance, such as flipping light switches or opening doors in the home.
-Based on these results, a first selection of tasks was selected for the Wizard of Oz design to present in the studies.
-In Section 4.1 the literature findings were used to group the responses from the online survey and to compare results related to the change in task prioritization over time in Section 5.1.
-2.3. Gaze as control input
+Such a system can achieve lower performance times and error rates when the user provides strategic intent and the autonomy supplies continuous, low-level stabilization, obstacle avoidance, and motion refinement (Farhadi et al., 2025).
+Advances in sensing, inference, modeling, and learning methods have led to novel developments in this research area (Selvaggio et al., 2021).
+Gaze can be used as an implicit or explicit control input for shared controls (Loke et al., 2025).
+As an implicit control input, natural gaze patterns are used to recognize or predict user intent.
+For example, gaze behaviors such as fixating on task-relevant objects before naming or manipulating them can be used and predicted before the desired task even begins (Admoni and Scassellati, 2017; Aronson and Admoni, 2020; Belardinelli, 2024; González-Díaz et al., 2024; Selvaggio et al., 2021).
+Huang and Mutlu demonstrated an example of such controls, in which ingredients were predicted by gaze patterns (Huang and Mutlu, 2016).
+Admoni et al.
+created a shared control system based on a POMDP (Partially Observable Markov Decision Process), which predicts single tasks from gaze patterns in an unstructured environment (Admoni and Srinivasa, 2016).
+Zhang et al.
+presented another approach based on the usage of a large language model (LLM) and vision foundation model (VFM), that recognizes user intent by interpreting labels created by the VFM and using the LLM to find common object combinations (Zhang et al., 2025).
+Wang used machine learning to recognize actions based on 3D gaze-related features to control a robot (Wang, 2021).
+Lastly, Farhadi et al.
+presented a review of BCI-related controls and shared control approaches (Farhadi et al., 2025).
+Challenges in predicting patterns of natural gaze can arise from multiple cognitive processes occurring in our brain at any time.
+As Belardinelli states, this issue worsens when performing activities more complex than grasping or selecting an object (Belardinelli, 2024).
+If incorrect or imprecise predictions are made, collisions or large penalties in completion time and user preferences can occur (Jain and Argall, 2019).
+As Cao et al.
+point out, timely and accurate estimation of intentions and, consequently, prediction remain major challenges (Cao et al., 2025).
+Gaze, as an explicit input, can vary in its function in control.
 The first eye-tracking controllers for robotic arms were developed as early as 2001 (Kim et al., 2001).
-Since then, several systems have been developed, of which only a few have become commercially available (Fischer-Janzen et al., 2024; Suryadarma et al., 2024; Brose et al., 2010).
-They differ in the control approach used to guide the robot.
-A popular design approach is directional control, where the robot moves in the direction of gaze (Alsharif, 2017; Sunny et al., 2021).
+Since then, several systems have been developed, but only a few have become commercially available (Fischer-Janzen et al., 2024; Suryadarma et al., 2024; Brose et al., 2010).
+One example of explicit control input is directional control, in which the robot moves in the direction of gaze (Alsharif, 2017; Sunny et al., 2021).
 Graphical User Interfaces (GUIs) are available in which the robot is moved with gaze-selected buttons, taking advantage of more accurate stationary eye-tracking systems (Dragomir et al., 2021; Stalljann et al., 2020).
-Going a step further, human intention in eye-tracking control can be predicted from task-related eye movements (anticipatory gaze (Aronson and Admoni, 2020)) based on the natural gaze behavior in hand-eye coordination (González-Díaz et al., 2024).
-Task anticipation has been explored to estimate task selection from the gaze pattern (Huang and Mutlu, 2016; Wang, 2021).
-It can be predicted before the action begins and during its execution (González-Díaz et al., 2024; Selvaggio et al., 2021).
-A crucial step in using gaze as control input is to decompose the behavior into subtasks to obtain specific gaze patterns that can be used to train AI models (Hayhoe and Ballard, 2014).
-Another application is to use gaze as a complement to generate more dense and accurate information, as is done in Brain-Computer-Interface development (Onose et al., 2012; Ivorra et al., 2018).
-Especially reading task intention from gaze allows to use low-level cognitive inputs to realize systems with high autonomy, making it easier to interact with the system (González-Díaz et al., 2024).
-In contrast to these possibilities, there are some challenges that need to be considered when implementing gaze as a control input:
-  • Unintended selection: When gaze is used as a pointing device, e.g.
-to select an object, rules must be applied to ensure, that not all fixations are selection triggers.
-This issue is called the Midas Touch problem and describes an unintentional selection of an object (Dünser et al., 2015; Jones et al., 2018; Stalljann et al., 2020).
-Reasons include unintentional and reflexive eye movements, overlapping objects, distraction, stress, and limited accuracy of the eye-tracking device (Holmqvist and Andersson, 2017).
-  • Context misinterpretation: It is caused by overlapping objects or task ambiguity.
-When objects are close together and the eye-tracker accuracy is not sufficient, the gaze can be tracked on the wrong object leading to faulty interactions between robot and object (Jones et al., 2018).
-In the interview conducted by Herr et al.
-(Herr et al., 2016) participants stated: “Well, now I look at you [the interviewer], but I actually focus on the bottle behind you.
+Such controls are easy to learn, provide a good sense of agency, and reduce cognitive load (Dragomir et al., 2021; Dziemian et al., 2016).
+However, controlling each of the robot movements separately is more time consuming.
+A popular approach, as outlined in Belardinelli’s work, is to use relevant objects in the scene and define them as Areas-Of-Interest (AOI).
+Fixations within these areas can be evaluated and, together with a certain dwell time, used as a selection method (Catalán et al., 2017; Zeng et al., 2017; Tostado et al., 2016).
+We refer to the work of Fischer-Janzen et al.
+and Suryadarma et al.
+for examples in realization, due to the number of related works (Fischer-Janzen et al., 2024; Suryadarma et al., 2024).
+A major limitation of these works is that they do not allow for the subsequent selection of multiple tasks, which is a major requirement for daily use, as demonstrated by Hagengruber et al.
+(Hagengruber et al., 2025).
+In this work, gaze is used as explicit input by focusing on the desired object.
+Challenges can occur in both implicit and explicit control inputs that need to be considered:
+  • Unintended selection: When the gaze is used as a pointing device, e.g., to select an object, rules must be applied to ensure that not all fixations do not trigger selections.
+These unintentional selections are described by the Midas Touch problem (Dünser et al., 2015; Jones et al., 2018; Stalljann et al., 2020).
+Reasons for this problem include unintentional and reflexive eye movements, overlapping objects, distraction, stress, and the limited accuracy of the eye-tracking device (Holmqvist and Andersson, 2017).
+When objects are close together and the eye tracker’s accuracy is insufficient, the gaze may be tracked to the wrong object (Jones et al., 2018).
+  • Context misinterpretation: The aforementioned parallel cognitive processes can result in task ambiguity.
+In an interview conducted by Herr et al.
+(Herr et al., 2016), participants stated: “Well, now I look at you [the interviewer], but I actually focus on the bottle behind you.
 That would be difficult for the robot to recognize”.
 This behavior suggests that attention to a particular location is sometimes unrelated to intention or cognitive process.
-  • Physiological factors: Physiologically, there are some factors that affect the accuracy of the gaze.
-Neurological conditions, fatigue, drooping eyelids and strabismus affect the functionality of eye-tracking devices (Holmqvist and Andersson, 2017).
-The robotic workspace and gaze movements differ in dimensions.
-Three-dimensional robot workspace has to be covered by two-dimensional eye movements.
-Eye tracking glasses have decreasing accuracy in the periphery of the field of view.
-When head movement is limited, as in the case of the stakeholder group, these peripheral regions are more likely to be used to fixate objects, leading to increased instances of contextual misinterpretation.
-When such challenges occur, it can lead to frustration and high cognitive load for the user to avert such situations (Dünser et al., 2015; Jones et al., 2018).
+  • Physiological factors: There are some physiological factors that affect the accuracy of the gaze.
+Neurological conditions, fatigue, drooping eyelids, and strabismus can affect the functionality of eye-tracking devices (Holmqvist and Andersson, 2017).
+Such challenges can lead to frustration and a high cognitive load for users trying to avoid them (Dünser et al., 2015; Jones et al., 2018).
+2.2. Characteristics in gaze-based shared controls
+Our system design was inspired by the impact factors on user acceptance of shared control found in the current literature.
+Many characteristics and technology acceptance models exist (Beer and Akanksha Prakash, 2011).
+Due to the multifaceted nature of user acceptance, we present only the factors that influenced our design.
+For more information on characteristics influencing user acceptance, see the works of Smarr et al., Rahman et al., Riek et al., and Beer et al.
+(Smarr et al., 2012; Rahman et al., 2025; Riek and Howard, 2014; Beer and Akanksha Prakash, 2011).
+We introduce three categories to facilitate an overview of the results: Technical challenges, level of autonomy, and user preferences.
+Technical challenges include ensuring the system’s reliability and functionality.
+The system must be robust in changing or unorganized environments, especially when the robot is mounted on an electric wheelchair (Selvaggio et al., 2021; Stanger et al., 1994).
+The system’s capabilities related to task performance also define its functionality in everyday life.
+We discussed such challenges in Section 2.1.
+The subtasks necessary to correctly complete a task must be understood and set in the proper context relative to each other.
+The level of autonomy is a subject of ongoing debate and can range from tele-operated systems to fully autonomous robots (Beer et al., 2014).
+Recent studies have shown that more autonomy does not necessarily align with user preferences (Nanavati et al., 2024b; Bhattacharjee et al., 2020).
+If a robot with a high level of autonomy produces more errors in task execution, users tend to reject the system (Bhattacharjee et al., 2020).
+This can lead to a loss of trust or a reduced sense of agency due to the system’s unpredictable behavior (Nanavati et al., 2024b; Baptista et al., 2025).
+Losey et al.
+summarized this as a duality: More automation reduces the user’s burden when performing tasks, but it also limits the mutual support when it ignores the user’s intent.
+The autonomous behavior reduces the operator’s workload when performing repetitive tasks (Selvaggio et al., 2021).
+As Loke et al.
+state, high mental workload contributes to the abandonment of assistive devices (Loke et al., 2025).
+Due to the potential shift in cognitive resources toward the task and away from precise robot movements, control can benefit from being easy to understand (Mazzola et al., 2025; Selvaggio et al., 2021).
+However, this may vary depending on the complexity of the task (Stanger et al., 1994; Shafti et al., 2019; Goldau et al., 2024), leading back to the initial question of what level of autonomy is most efficient and reliable in assisting the user.
+Users’ preferences vary based on personal metrics such as the user’s age, level of disability, living arrangements, and familiarity with technology (Stanger et al., 1994; Smarr et al., 2012; Nanavati et al., 2024b).
+Incorporating user preferences influences perception of and satisfaction with the system (Cakmak et al., 2011; Canal et al., 2021).
+Parameters such as speed can influence the user perception (Bhattacharjee et al., 2020).
+Another example is the participant’s preference for handing over an object with a particular orientation (Cakmak et al., 2011).
+The ability to adapt the system seamlessly to users’ preferences regarding the level of assistance has given rise to shared autonomy approaches.
+Those are based on the autonomy’s understanding of the user’s intentions and of the surrounding environment (Selvaggio et al., 2021).
+In this context, research focuses on arbitration strategies that adapt the level of autonomy to the user’s preferences, as demonstrated in the works of Dragan et al., Losey et al., and Jain et al.
+(Dragan and Srinivasa, 2013; Jain and Argall, 2019; Losey et al., 2018).
+The arbitration scheme, together with the interaction paradigm, describes the collaboration between human and robot (Farhadi et al., 2025).
+When designing well-balanced shared control, a trade-off must be found for all technical and ethical factors. Section 3.1 presents a solution that considers and balances the different influences.
+2.3. Tasks in assistive robot manipulation
+Edemekong et al.
+describe Activities of Daily Living [(ADLs)] as basic routine tasks that most healthy individuals can perform without assistance (Edemekong et al., 2025).
+ADLs are often used to evaluate a person’s level of disability, as in the WHODAS 2.0 (World Health Organization, 2010).
+Nanavati et al.
+used ADLs in their review to compare the responses of individuals requiring ADL assistance with the work of researchers in the field of assistive robotics.
+A major misalignment was found.
+While most researchers focus on mobility and eating, people in need of assistance state that they need help with going out, managing money, doing housework, preparing meals, bathing, and dressing (Nanavati et al., 2024b).
+Petrich et al.
+took a different approach, examining the behavior of people in everyday life and uncovering results quite different from those previously mentioned.
+In their study, tasks such as reaching to pick up and place items, opening and closing doors, and using switches and buttons were most important, according to video analysis.
+These results differed from the participants’ survey answers (Petrich et al., 2022).
+These results highlight that different approaches can reveal various aspects of task prioritization and preferences in HRI.
+Smarr et al.
+explored robot acceptance among the elderly and found that the elderly accept robots more readily when they assist with learning new skills and hobbies.
+However, human support was preferred for communication, which emphasizes the importance of social contact (Smarr et al., 2012).
+The examples presented above show that assistive robotic arms cannot directly assist with activities such as bathing or managing finances.
+It is important to distinguish between robotic arm-related tasks and others.
+As stated by Hagengruber et al., the purpose of assistive robotics is to carry out various tasks in succession in order to depict a realistic daily routine (Hagengruber et al., 2025).
+They presented an sEMG-based control system that could perform a drinking task, including retrieving a cup from a drawer and pouring a drink.
+Chi et al.
+demonstrated that dividing complex tasks into subtasks facilitates accomplishment.
+For example, they divided an eating task into subtasks such as getting a spoon, scooping food, and bringing it near the user’s mouth (Chi et al., 2025).
+Using gaze as a control input, one can obtain gaze patterns specific to each subtask that can be used to train AI models (Hayhoe and Ballard, 2014).
+Robots can perform daily tasks, such as household chores and personal hygiene, as well as medical procedures (Herr et al., 2016; Rahman et al., 2025; Bien et al., 2003; Rahman et al., 2025; Wang et al., 2020).
+The preferred tasks vary depending on different influences, such as the user’s attitude towards a task.
+Delegating household tasks to robots is one example of this (Herr et al., 2016).
+Similarly, Petrich et al.
+suggested automating tasks that are not prioritized by the community to reduce cognitive load (Petrich et al., 2022).
+Lastly, one must determine whether a robotic arm is the right tool for the task at hand.
+For example, Petrich et al.
+stated that robots can help with using electronics, such as smartphones; however, it is more convenient to use a hands-free phone (Petrich et al., 2022).
+Similar results were found for toileting, where people preferred the aid of caregivers and smart home applications.
+Hagengruber et al.’s work, in particular, demonstrates the advantage of using multiple assistive devices together.
+For example, it simplified opening a drawer by holding the handle with the robot and moving the wheelchair backward (Hagengruber et al., 2025).
+In addition to the work of Nanavati et al.
+(Nanavati et al., 2024b), our review revealed that the four most frequently mentioned tasks, suitable for assistive robots, were food related tasks (serving a meal, eating) (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003; Canal et al., 2021; Bhattacharjee et al., 2020; Nanavati et al., 2024a), drinking related tasks (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003), picking up objects (Bien et al., 2003; Stanger et al., 1994; Jardón et al., 2012; Chang and Park, 2003; Nanavati et al., 2024b), and personal hygiene (Stanger et al., 1994; Jardón et al., 2012; Nanavati et al., 2024b).
+Personal hygiene included various tasks such as shaving (Bien et al., 2003; Chang and Park, 2003), washing the face and hands, brushing teeth, combing hair, shaving, applying makeup (Jardón et al., 2012), and dressing (Canal et al., 2021; Nanavati et al., 2024b).
+The importance of being able to manipulate a wide variety of objects benefits the ability to get around, e.g., with an electric wheelchair, by being able to move blocking objects out of the way (Nanavati et al., 2024b).
+Other miscellaneous tasks were mentioned with taking medicine (Nanavati et al., 2024b), working (Nanavati et al., 2024b), leisure activities (Stanger et al., 1994; Nanavati et al., 2024a), turning on and off switches (Bien et al., 2003; Chang and Park, 2003), opening and closing doors (Bien et al., 2003; Brose et al., 2010; Chang and Park, 2003), scratching oneself (Bien et al., 2003; Chang and Park, 2003), and making tea.
+Sources from 2003 show that changing CDs and removing paper from fax machines are examples of how task preferences change over time (Bien et al., 2003; Chang and Park, 2003).
+As the accessibility of smart home applications continues to evolve, it is expected that more tasks will become less important for robot assistance, such as flipping light switches or opening doors.
+Based on these results, tasks were selected for the Wizard of Oz design.
+In Section 4.1, the literature findings were used to group the responses from the online survey.
+In Section 5.1, they were used to compare the results in terms of changes in task prioritization over time.
 2.4. Participation in assistive robot design
-Aktion Mensch, a german organization for social projects, states that current assistive devices often fail to meet the needs of users (Borgstedt and Möller-Slawinski, 2023).
-In eye-tracking controls most studies recruit non-disabled people as participants (Fischer-Janzen et al., 2024).
-Reasons why people with disabilities are rarely included in studies are recruitment challenges, small sample sizes, and transportation logistics (Nanavati et al., 2024a).
-As shown by Stalljann et al., control inputs are experienced differently by disabled and non-disabled participants (Stalljann et al., 2020).
-Including people with disabilities helps in the design process of assistive devices by gaining insight into their lives and needs (Baumeister et al., 2021; Stanger et al., 1994; Nanavati et al., 2024a).
+The previous section presented characteristics that influence user acceptance and robot abandonment.
+Several authors argue that involving the community and people with disabilities in the design of assistive devices is more advantageous than involving people without disabilities.
+Including PSMD in the design process of assistive devices helps designers gain insight into the lives and needs (Baumeister et al., 2021; Stanger et al., 1994; Nanavati et al., 2024a).
+Arboleda et al.
+point out that end users’ perspectives can differ from those of other groups, such as caregivers (Arevalo Arboleda et al., 2021).
+As Stalljann et al.
+demonstrate, control inputs are experienced differently by disabled and non-disabled participants (Stalljann et al., 2020).
+Most studies on eye-tracking controls recruit non-disabled people as participants (Fischer-Janzen et al., 2024).
+Reasons PSMD are rarely included in studies include recruitment challenges, small sample sizes, and transportation logistics (Nanavati et al., 2024a).
 Nanavati et al.
-(Nanavati et al., 2024a) discusses approaches based on three-dimensional key features to find the most appropriate study design for the goal of the study.
-The first dimension decides between individual or community-level participation.
-With individuals, the chance of a deeper insight into the user’s life and preferences can be gained, while community work can lead to a broad insight from many participants.
-It also affects the time needed to conduct the studies and the connection between participant and experimenter.
-The second dimension highlights the logistical burden on participants and researchers.
-People with severe physical disabilities are usually dependent on their caregiver.
-Depending on the robot setup, it is easier to conduct studies at the participant’s location, to reduce the need of special transportation and availability of caregivers.
+(Nanavati et al., 2024a) discuss approaches based on three-dimensional key features to determine suitable study designs for achieving the study’s objective.
+The first dimension determines whether participation is at the individual or community level.
+Individual participation provides the opportunity for deeper insight into the user’s life and preferences, while community work provides insight from many participants.
+This dimension also affects the time required to conduct the study and the connection between participants and experimenters.
+The second dimension highlights the logistical effort required of participants and researchers.
+PSMD participants may require assistance with transportation to study sites, which can facilitate conducting studies at the participants’ locations.
+Arboleda et al.
+stated that conducting the study itself might disrupt participants’ daily routines, which could induce stress (Arevalo Arboleda et al., 2021).
+Therefore, other possibilities can be considered, such as online interviews.
+Depending on the robot setup, it may be easier to conduct studies at the participant’s location to reduce the need for special transportation and availability of medical professionals.
 The third dimension describes the benefits to researchers and the community.
-Participants have an intrinsic motivation to participate in the design process of assistive devices and robots.
-They can benefit by providing impetus to adapt the systems to their needs.
-Researchers benefit by potentially making important connections for future work.
-Although it does affect the length of the study, it is worthwhile to involve people with disabilities in the design process. These findings influenced the study design in Section 3.4.
+PSMD, their families, and medical professionals are intrinsically motivated to participate in the design process of assistive devices and robots.
+They can benefit and contribute by providing insight to adapt the systems to their needs.
+These findings demonstrate the value of involving PSMD in the design process and influenced the study design in Section 3.3.
 3. Methods
-In the introduction and background, we presented influencing factors and user challenges that affect the adoption rates of robotic arm systems.
-These challenges and factors were partially incorporated into the design of the system, presented in Section 3.1.
-Several groups of interest were interviewed with two study designs, to obtain broader feedback on the shared control design.
-The setup described in Section 3.3 and 3.2 was used in both studies.
-An online survey was sent to stakeholder groups, as well as their families and medical professionals and is presented in Section 3.4.
-Measurements were conducted to verify technical functionality and the reaction of participants interacting with the robot in person (see Section 3.5).
+In Sections 1 and 2, we presented the influencing factors and user challenges that affect acceptance of eye-tracking-driven shared controls.
+These factors and challenges were incorporated into the system design presented in Section 3.1.
+Several groups of interest were interviewed with two study designs, to obtain broader feedback.
+The setup described in Section 3.2 was used in both studies.
+An online survey was sent to stakeholder groups, their families and medical professionals.
+The results are presented in Section 3.3.
+Measurements were conducted to verify technical functionality and participants’ perception of the system interacting directly with the robot (see Section 3.4).
 3.1. Shared control design choices
-The initial idea of the system is to include a shared control in which parts of the robot movement are automated, such as trajectory planning, grasping and object manipulation.
-The level of autonomy is presented following the guidelines of Beer at a.
+The initial idea for the system is to include a shared control, which automates parts of the robot movement, such as trajectory planning, grasping, and object manipulation.
+The level of autonomy is presented according to the guidelines of Beer et al.
 (2014) (Beer et al., 2014).
+The following considerations were used to create the Wizard of Oz design and describe the interaction to the participants.
 What task is the robot to perform?
-This shared control is designed to assist people with severe physical disabilities in everyday tasks.
-Ongoing work is realizing a framework that can be extended to include everyday activities previously unknown to the robot.
-The first approach includes the tasks of grasping and handing drinks, filling a glass with a drink, and grasping and placing objects in locations selected by the user.
-These tasks can be performed in unsorted environments such as at home and in public places, e.g.
-a restaurant.
-These tasks represent two of the most preferred tasks for robotic assistance in everyday life, as shown in Section 2.2.
+This shared control system is designed to assist PSMD with daily tasks.
+Ongoing work is developing a framework that can be expanded to include new ADL, thereby increasing the range of tasks it can perform.
+The first approach includes the tasks of picking up food with a fork, turning on a light switch, picking up and placing a block, scooping from a bowl, pouring water into a glass, and handing it to the participant.
+These tasks have to be performable in unstructured environments, such as at home and in public places like restaurants.
+As shown in Section 2.3, these tasks are preferred for robotic assistance in everyday life.
 What aspects of the task should the robot perform?
-The chosen input modality is eye-tracking, since eye movement is often unrestricted or only slightly restricted in contrast to other motor functions such as arm movement in cases of locked-in syndrome or cerebral palsy.
-The control is realized by fixating the desired object and the robot decides based on the selected object which available task to perform.
-Multiple objects can be selected by gaze to make task selection more robust.
-For example, looking at a cup could indicate that the user wants to drink from it, move it, fill it, or combine available tasks.
-By adding the ability to combine the cup with other objects, such as selecting a bottle or a location, the task can be distinguished into refilling the cup or placing it in the selected location.
-Based on object and ask information, the robot performs the task autonomously. It determines the location of the object, a suitable grasp related to the task, and calculates the optimal trajectory.
+The chosen input modality is eye-tracking, because eye movement is often unrestricted or only slightly restricted in cases of locked-in syndrome or cerebral palsy, in contrast to other motor functions, such as arm movement.
+The robot is controlled by fixating the desired object, and the robot decides which available task to perform based on the selected object.
+Selecting multiple objects with the gaze makes task selection more robust.
+For instance, looking at a cup could indicate that the user wants to drink from it, move it, fill it, or combine available tasks.
+Adding the ability to combine the cup with other objects, such as a bottle or a location, distinguishes the task as refilling the cup or placing it in the selected location.
+Based on the information about the object and the task, the robot performs the task autonomously.
+The robot determines the object’s location, selects a suitable grasp related to the task, and calculates the optimal trajectory.
 To what extent can the robot perform those aspects of the task?
-The user is always in control of the robot.
-Emergency stop methods are considered to ensure the well-being of the user.
-Validation methods are planned to intervene in case of incorrect task execution before the robot starts the task.
-Settings are planned to reduce the robot’s velocity if desired.
-With the implementation of dwell time selection of objects, incorrect interaction with the object is reduced with natural gaze behavior, thus minimizing the Midas Touch problem.
-What level can the robot’s autonomy be categorized? According to the classification of Beer et al. (Beer et al., 2014) robot autonomy is categorized as Shared Control with Human Initiative.
+The user is always in control of the robot in order to maintain a sense of agency (discussed in (Nanavati et al., 2024b)).
+Validation methods are planned to intervene in case of incorrect task execution before the robot begins (suggested in (Baptista et al., 2025)).
+Settings are planned to reduce the robot’s velocity, if desired (suggested by (Bhattacharjee et al., 2020)).
+Implementing dwell time for selecting objects has the potential to reduce incorrect interactions with objects through natural gaze behavior, thus minimizing the Midas Touch problem.
+What level can the robot’s autonomy be categorized? According to the classification of Beer et al. (Beer et al., 2014), robot autonomy is categorized as Shared Control with Human Initiative.
 How might autonomy influence HRI variables?
-Reliability and functionality of the system must be given.
-Safety for the user must be ensured and is part of ongoing discussions.
-This system is designed to be easy to learn and use.
-Since there is no physical interface such as a display, the correct selection of the object depends on the accuracy of the eye tracker.
-Objects are selected by focusing on them in real-world settings.
-If the eye tracker is incorrectly calibrated, the gaze may not be tracked to the object due to high accuracy errors.
-This raises the question of how serious this error is in this setup, which will be tested with the Wizard of Oz setup.
-A wearable eye-tracker is used as input device.
-Influences on social variables such as aesthetics or the occupation of the voice by voice control, which could lead to a rejection of the system, are minimized.
+The system must be reliable and functional.
+User safety is part of ongoing discussions and must be ensured.
+The system is designed to be easy to learn and use.
+In real-world settings, objects are selected by focusing on them.
+However, the accuracy of the eye tracker limits the correct selection of the desired object when gaze is tracked outside the area of interest (AoI).
+This raises the question of how accuracy errors influence the functionality of this approach.
+This will be tested using the Wizard of Oz setup.
 3.2. Wizard of Oz design
-Based on the results of the literature review, we decided to do an early evaluation of the design if it meets the expectations for the influencing factors given by the community.
-This approach has the advantage of shortening the development time by intervening in case of low acceptance or functional challenges.
-Therefore, the in the following used system represents a Wizard of Oz design, where the autonomously controlled robot tasks are performed by the experimenter, as seen in Figure 1.
-The objects were placed in a predefined position in the robot’s workspace.
-All trajectories needed to realize the automated task execution were hard-coded and could be selected by the experimenter clicking on the program.
-The safety of the participant was ensured by separating the workspace of the participant and robot.
-The experimenter was informed about the task sequence of the user and observed the participant’s visualized eye movements in a live video on the monitor.
-As soon as the participant fixated on an object, the corresponding task was started via the robot’s web interface.
-Five tasks were selectable with the Wizard of Oz system: Picking up food with a fork, turning on a light switch, picking up and placing a block, scooping from a bowl, and filling a glass with water and handing it to the participant.
-Each task was related to different objects, as shown in Figure 2.
-The selection were based on the initial idea for the system and the outcomes of the literature review.
-A video of task performance was created to be implemented in the online survey to visualize the function of the system (Section 3.4).
+An early system evaluation was conducted to determine whether the design meets the community’s expectations for the found influencing factors.
+This approach shortens development time by intervening early in the event of low acceptance or functional challenges.
+In the Wizard of Oz design, the experimenter performs the autonomously controlled robot tasks, as seen in Figure 1.
+Table 1 presents the Wizard of Oz system according to the guidelines presented by Riek (Riek, 2012).
+Figure 2 shows the five selectable tasks with the Wizard of Oz system.
+A video of task performance was created and implemented in the online survey to demonstrate the system’s functionality (Section 3.3).
 This video is included in the Supplementary Materials.
@@
 Both studies were reviewed and approved by the ethics committee of the first author’s institution.
+Table 1. Wizard of Oz presentation after Riek (Riek, 2012) stating experimental components to the robot, user, wizard and to general information.
+Experimental Component Online survey Hands-on study
+Robot 7-DoF Kinova Gen3
+Level of autonomy Teleoperation
+Robot capabilities Food item, scooping food, filling cup Food item, pick and place, filling cup, light switch
+No. of participants 34 24
+Demographics Age: $40\pm 27$; Age: $30.5\pm 8$;
+14 PSMD, 9 family members, 11 medical professionals; 11 non-prior experience group, 13 prior-experience group;
+14 male, 17 female, 3 diverse 7 female,17 male
+Instructions See Section 3.3.3 See Section 3.4.1
+Convincing simulation Not asked Yes
+Wizard Experimenter took the role of the wizard
+Familiar with hypothesis? Yes
+Production variables - See Section 3.4.1
+Recognition variables - See Section 3.4.1
+Wizard error Video representation, no variation between tests Experiments were repeated when robot malfunctioned, occurrence: 1.
+Location Unknown, not asked due to privacy regulations At the laboratory
+Environmental constants Unknown Setup as shown in Figure 2
+Scenario As shown in video (see Supplementary Materials) See additional information to scenario.
+Iterative design process No No
+Limitations See Section 6
+Reason for omitting the item ”Experimental Component” from the Wizard presentation of Riek et al.: The hypothesis for the robot was to reproduce each participant’s trajectory and create as little variation as possible.
+Thus, the robot’s software computed the trajectory planning.
+The points between the trajectories were programmed so that the overall task trajectory would be reproducible between participants.
+As Arboleda et al.
+indicated, robot behavior is perceived differently in virtual and real presentations (Arevalo Arboleda et al., 2021).
+Therefore, a crucial aspect of the behavioral hypothesis was to compare the two presentations.
+Distinct reactions of the participants were not previously hypothesized, since the goal of this work was to gain insights into user expectations and behavior.
+Additional information on the item ”Scenario”: The objects were placed in predefined positions within the robot’s workspace.
+All of the trajectories necessary for the automated task execution were hard-coded and could be selected by the experimenter by clicking on the program.
+The participant’s safety was ensured by separating their workspace from the robot’s.
 Figure 1.
@@
 The home position is used as indication that the participant can select the next task.
-[…]
 Wizard of Oz design: The participant selects an object by focusing on it with his or her gaze.
@@
 Figure 2. Nine interactable objects were presented with which five tasks were performed. Participants were shielded from visual distractions by green walls surrounding the experimental area.
-[…]
 The figure shows the setup with nine objects placed either on the table or, in the case of the light switch, on a plywood wall that could be selected to perform five tasks.
@@
 The emergency stop button was placed on the table in the front left corner in front of the light switch so that participants could reach it if they felt unsafe.
-3.3. Technical setup
-The Wizard of Oz setup consists of a robot (Kinova, Gen3 with a gripper Robotiq, 2F-85) controlled via the standard web interface provided by the manufacturer.
+3.3. Online survey
+The online survey was chosen as the data collection tool because it provided a standardized way to present information and collect feedback.
+The goal was to obtain comprehensive feedback from PSMD, their family members, and medical professionals in order to evaluate the system’s usability and accessibility.
+Providing an overview of the system to the stakeholder group via online survey and video minimized the logistical challenges described in Section 2.4.
+This allowed immobile individuals to participate using smart devices.
+Unlike a real-world demonstration, where the authors could only visit a few facilities, people living at home could participate.
+Additionally, the survey could be completed at the participant’s preferred time, which was convenient for those with complex schedules or who required assistance.
+Due to the anticipated limited availability of accessible computer controls, an extension was designed for family members and medical professionals to obtain more comprehensive feedback.
+The purpose of the online survey was to answer the following questions:
+  • Are potential users able to use the system?
+  • Which tasks should be prioritized to maximize the robot’s usefulness in everyday life?
+  • Were the participants satisfied with the presented system?
+3.3.1. Survey development and structure
+The survey focused on investigating five topics, which are listed in Table 2.
+First, questions were asked about the participants’ demographics to provide comparable and reusable baseline data.
+After the five topics were covered, participants were given the opportunity to provide feedback on the presented system and make general remarks.
+As stated in Section 2.4, perceptions of assistive devices can differ between end users and individuals in close contact with them, such as medical professionals and family members.
+Our goal was to receive a broad perspective on the prototype, which is why the survey varied between these three groups.
+Questions about the support situation were designed to evaluate the survey’s accessibility.
+Poor accessibility can be indicated by how Q4(PSMD) is answered.
+A higher number of relatives and medical professionals who did not involve the person they care for can indicate time constraints.
+Furthermore, if everyone in Q4(relatives/friends) had stated that the person they care for had answered the survey, the answers would need to be weighted due to potential bias.
+The topics of disabilities and assistive devices focus on whether the respondent is the intended audience.
+We implemented questions to reduce responses from people who misidentified their group affiliation.
+The topics also focus on whether the respondent is familiar with assistive devices, including robotic arms.
+The literature contains cases in which users reacted differently to devices when they had prior knowledge.
+This may affect their reaction to the presented system.
+The questions regarding the robot presentation were developed to identify signs of abandonment.
+Robotic arms are not widespread.
+Upon inquiry, two companies that sell assistive robotic arms approved as medical devices reported selling around 1,200 models in Germany, Switzerland, and Austria.
+Since this strongly affects the system’s perception, participants were asked if they had experience with robotic arms.
+Lastly, task preferences were requested.
+As technology continues to evolve, the aim was to update the ongoing literature.
+As presented in Section 2.3, task preferences change quickly with the advent of new technology, such as smart home applications and hands-free smartphones.
+Some tasks, like using a fax machine, are no longer relevant.
+We asked participants to rank the importance of three tasks against each other.
+Throughout the survey, we chose open-ended questions to receive rationale for the participants’ answers to yes/no questions, such as reasoning for being unable to use the system.
+After initial development, the survey was discussed and refined.
+Table 2. Breakdown of topics asked in the survey and related questions.
+Question topic PSMD Rel./Fr. Med. prof.
+Support situation Q4+Q5 Q4+Q5 Q4
+Disability & assistive devices Q6-Q8 Q6+Q7 -
+Robotic arm use Q9-Q12 Q9+Q10 Q9+Q10
+WoZ satisfaction Q13+Q14 Q11+Q12+Q14+Q15 Q11+Q12+Q14+Q15
+Task preferences Q15 Q13 Q5+Q13
+As a result, the online survey consisted of 31 questions, divided into three groups, as well as a video describing the eye-tracking control of the robotic arm, structured as shown in Figure 3.
+In addition to the tasks, the video showed the calibration process.
+The exact questions asked are provided in Appendix A.
+All survey questions, along with the participants’ responses, are available in the Supplementary Materials.
+The survey was designed using LimeSurvey.
+Figure 3. Distribution of questions asked for each group. Q outlines the context of the questions and the number of the questions for each category is given.
+Distribution of questions asked for each group. Q outlines the context of questions asked and the number of the questions for each category is given.
+3.3.2. Distribution method
+The survey was available online from September 18, 2023, to January 22, 2024.
+32 organizations in Germany, Switzerland, and Austria were emailed to reach the stakeholder groups.
+Seven organizations, including FGQ (Paraplegic Support Association), LIS e.V.
+(Locked-In Syndrome Association), Diakonie (social services), Lebenshilfe (nonprofit association), and the Swiss Paraplegic Centre, shared the survey.
+The authors’ private accounts and their institutions shared the survey a total of three times on social media.
+FGQ shared it in their private Facebook group, and the Swiss Paraplegic Centre posted about it in their private forum.
+3.3.3. Participant instructions
+First, participants were informed about the content and purpose of the study.
+Then, they were asked to give their consent to participate.
+Participants were then guided through the survey.
+Participants could skip questions at any time if they felt uncomfortable answering them.
+The video played when clicked on, and could be repeated if necessary.
+To offer different resolutions, participants could also open the video on YouTube.
+3.3.4. Evaluation method
+A frequency analysis was performed to evaluate users’ perceptions of the system and their desired tasks.
+As introduced in Section 2.3, the tasks were categorized as follows: eating, drinking, picking up objects, environmental interactions, personal hygiene, and miscellaneous.
+3.4. Hands-on user study
+In addition to the stakeholder evaluation of the system, a Wizard of Oz design was employed to analyze real-world interactions.
+Evidence from the online survey suggests that participants can use the system based on their answers.
+However, challenges affecting the system’s functionality may occur due to the eye tracker’s insufficient accuracy or natural gaze behavior.
+Participants interacted with the robot to record gaze data, which was then used to evaluate hardware constraints.
+This study addresses the following questions:
+  • Would the prototype eye-tracking controller work in the controlled environment of the lab?
+  • What technical difficulties arise from the design?
+  • Does perception vary between the hands-on study and the online survey?
+3.4.1. Procedure and participants
+To minimize the influence of learning effects or behavioral adaptation, 24 participants were involved, creating 24 emerging permutations in the task sequence.
+Each participant provided informed consent after the procedure was explained.
+They were asked to sit in a chair in front of the robot workspace.
+The eye-tracker was set up and calibrated as described in (Pro, 2024).
+Participants started the task by fixating on one of two objects in the drinking, eating, light-switching, and picking-and-placing tasks with gaze (see Figure 1).
+Participants chose the objects according to their preference.
+The experimenter performed the role of the automation, as described in Section 3.2.
+Production and recognition variables were discussed prior to the test.
+Within the first second of selection, the wizard selected the corresponding task via the web interface.
+The robot’s reaction took between one and five seconds from the time the first reaction was recognized, including the wizard’s reaction time.
+As shown in Table 1, one wizard error occurred, causing the robot to stop shortly after the experiment started due to an incorrect mode selection.
+This experiment’s results were discarded, and it was repeated with another participant to ensure naivety to the procedure and function of the Wizard of Oz setup.
+Wizard training: The experimenter was informed of the user’s task sequence and observed the participant’s eye movements in real time on the monitor.
+As soon as the participant fixated on an object, the corresponding task started via the robot’s web interface.
+3.4.2. Data analysis
 Data were recorded using wearable eye-tracking glasses (Tobii, Tobii Pro Glasses 3, accuracy = 0.6).
 Gaze visualization, data analysis, and post-processing were performed using Tobii Pro Lab (Pro, 2014).
-The control computer to move the robot and visualize the participant’s gaze was placed 2 m behind the participant in the hands-on study so that the user was unaware of the experimenter’s role.
-3.4. Online survey
-The online survey was chosen as the data collection tool because it provided a standardized way to present information and collect feedback.
-The purpose is to gain broad feedback from people with disabilities, their family members and caregivers to assess the usability and accessibility of the system.
-The logistical challenges described by Nanavati et al.
-(Nanavati et al., 2024a) were minimized by providing an overview of the systems to the stakeholder group via online survey and video.
-It was advantageous because it allowed immobile people to participate through the use of smart devices.
-People living at home could participate, as opposed to a real-world demonstration where the authors could only visit a few facilities.
-In addition, it offered the convenience of completing the survey at the participant’s preferred time, accommodating those with complex schedules or who require assistance.
-Due to the anticipated limited availability of accessible computer controls, an extension was designed for family members and caregivers to obtain more comprehensive feedback.
-The purpose of this online survey was to answer the following questions:
-  • Are potential users able to use the system?
-  • Which tasks are preferred so that the robot has the greatest benefit in everyday life?
-  • Were the participants satisfied with the presented system?
-3.4.1. Distribution method
-The survey was available online from September 18, 2023 to January 22, 2024.
-32 organizations in Germany, Switzerland, and Austria were contacted by email to reach the stakeholder groups.
-With reference to the distribution channels mentioned by Nanavati et al., we selected the following organizations and platforms (Nanavati et al., 2024a).
-Seven organizations shared the survey, such as FGQ (Paraplegic Support Association), LIS e.V.
-(Locked-In Syndrome Association), Diakonie (social services), Lebenshilfe (non-profit association), and the Swiss Paraplegic Centre.
-On social media, the survey was shared a total of three times by the authors’ private accounts and by their institutions.
-FGQ shared it in their private Facebook group and a post was made in the private forum of the Swiss Paraplegic Centre.
-Figure 3. Distribution of questions asked for each group. Q outlines the questions’ context and the number of the questions for each category is given.
-[…]
-Distribution of questions asked for each group. Q outlines the context of questions asked and the number of the questions for each category is given.
-3.4.2. Survey structure
-The online survey consisted of 31 questions divided into three groups and a video describing the eye-tracking control of the robotic arm structured as shown in Figure 3.
-The video showed the calibration process, the light switch task, the pick and place task, and the scoop task.
-In the survey, people with disabilities were asked questions to assess their level of disability and daily performance.
-Their need of the control modality, their access to assistive devices, and their evaluation of the presented eye-tracking control was surveyed.
-Open-ended questions were used to identify tasks that the robot should be able to perform.
-Family members and caregivers were asked about the amount of time they spend providing care and their estimation of the number of people with physical disabilities in their care facility.
-In contrast to the questions above, family members and caregivers were first asked what tasks they would like a robot to perform in general, regardless of the type of robot.
-In this situation, the participants were unaware of the robotic system.
-Then the system was presented and questions similar to the first questionnaire were asked.
-Finally, all groups were given the opportunity to provide feedback on the survey and the system itself.
-All of the questions asked in the questionnaire, along with the participants’ responses, are available in the Supplementary Materials. The survey was designed with LimeSurvey.
-3.4.3. Evaluation method
-A frequency analysis was performed to evaluate the users’ perception of the system and the desired tasks.
-The tasks were categorized as introduced in Section 2.2: tasks related to eating, drinking, picking up objects, getting along, personal hygiene, and miscellaneous tasks.
-3.5. Hands-on user study
-In addition to the stakeholder evaluation of the system, the Wizard of Oz design was used to analyze real-world interaction.
-Evidence from the online survey suggests that participants are able to use the system based on their motor function and understanding of the system.
-The system realization might pose challenges in terms of functionality due to insufficient accuracy or natural gaze behavior.
-Participants were asked to interact with the robot to record gaze data that would be used to evaluate hardware constraints.
-Questions answered by this study include:
-  • Would the prototype eye-tracking controller work in the controlled environment of the lab?
-  • What are the technical difficulties with the design?
-  • Does participant perception vary greatly between the hands-on study and the online survey?
-3.5.1. Procedure and participants
-24 participants took part in the study (7 female/17 male, mean age = 30.5, 11 non-prior experience group/13 prior-experience group with robots) to minimize the influence of learning effects or behavioral adaptation related to the 24 emerging permutations in the task sequence.
-Each participant was informed and gave consent to be included.
-They were asked to sit in a chair in front of the robot workspace.
-The setup and calibration of the eye-tracker was performed as described in (Pro, 2024).
-Task start was selected by looking at one of two objects in the drinking, eating, and picking and placing tasks (see Figure 1).
-Participants chose the objects according to their preference.
-The experimenter performed the role of the automation as described in Section 3.2.
-3.5.2. Data analysis
-An evaluation of the Areas of Interest (AoI) and the dwell time is necessary to estimate the need for error minimization methods and to determine constraints for the object recognition algorithms.
-In combination with this technical evaluation of the system, the perception of the non-disabled participants was analyzed.
+In the hands-on study, the computer controlling the robot and visualizing the participant’s gaze was placed 2 m behind the participant so that the user was unaware of the experimenter’s role.
+Evaluating the areas of interest (AoI) and dwell time is necessary to estimate the need for error minimization methods and determine constraints for object recognition algorithms.
+Along with this technical evaluation of the system, we analyzed the perception of the non-disabled participants.
 In the Tobii Pro Lab events, Times of Interest (ToI), gaze movements during object selection, and Areas of Interest (AoI) were defined and labeled on the data.
-Metrics such as dwell times, which describe the total fixation time of an AoI relative to the ToI, and the number of missed hits were calculated by Tobii Pro Lab and exported for further analysis.
-Based on these parameters, the technical evaluation was performed.
-Insights into the factors mentioned in Section 2.3 and Section 2.1 are gained by interpreting gaze data through Times of Interest (ToI), type of gaze movement, and Areas of Interest (AoI) to distinguish dwell times of AoI hits.
-These parameters can be used to describe the scene based on visual attention.
+Tobii Pro Lab calculated metrics such as dwell times, which describe the total fixation time of an AoI relative to the ToI and the number of missed hits, and exported them for further analysis.
+The technical evaluation was performed based on these parameters.
+Interpreting gaze data provides insights into the factors mentioned in Section 2.1.
+ToI, AoI, and dwell time can be used to describe the scene based on visual attention.
 In order to facilitate the understanding of such terms, definitions according to Holmqvist and Nyström (Holmqvist and Andersson, 2017) and Tobii (Pro, 2014) are presented below.
-  • Event: Describes in Tobii Pro Lab the time stamp when an event occurs. In this case it marks the start and end time of a task.
-  • Time of Interest (ToI): Interval between two related events
-  • Area of Interest (AoI): A polygon corresponding to the outline of an object. The resulting area describes a visual stimulus.
-  • AoI hit: Describes the entry of the tracked gaze into an AoI.
-  • Dwell Time: Describes the time the gaze remains within an AoI after the initial entry.
-In this study, the gaze was allowed to exit and enter the AoI.
-It was the sum of time while the gaze stayed within an AoI.
-  • Type of eye movement: In this study, fixations and saccades were of most interest.
-Fixations refer to the time the eyes remain stable and focused on a specific point or area of interest.
-It was not distinguished from smooth pursuit, in which the head is moved while the gaze rests on a fixed point in the scene.
+  • Event: In Tobii Pro Lab, describes the timestamp when an event occurs. In this case, it marks the start and end times of a task.
+  • Time of Interest (ToI): The interval between two related events.
+  • Area of Interest (AoI): A polygon that corresponds to the outline of an object. The resulting area describes a visual stimulus.
+  • AoI hit: The entry of the tracked gaze into an AoI.
+  • Dwell Time: The time that the gaze remains within an AoI after the initial entry.
+In this study, the gaze was allowed to exit and re-enter the AoI.
+Dwell time is the sum of time while the gaze stays within an AoI.
+  • Type of eye movement: In this study, fixations and saccades were of primary interest.
+Fixations refer to the time that the eyes remain stable and focused on a specific point or area of interest.
+They were not distinguished from smooth pursuit, which occurs when the head moves while the gaze remains fixed on a point in the scene.
 Saccades are rapid changes in gaze direction.
 They connect fixations.
-Other eye movements such as vergence movements and microsaccades were not investigated.
-For further information on gaze movement we refer to Holmqvist and Nyström (Holmqvist and Andersson, 2017).
+Other eye movements, such as vergence movements and microsaccades, were not investigated.
+Further information on gaze movement can be found in Holmqvist and Nyström (Holmqvist and Andersson, 2017).
 4. Results
 4.1. Online survey
-20 out of 71 participants completed the survey.
-Of the 51 participants who filled out the survey, 36 surveys with no responses (N/A) had to be excluded.
-If participants completed part of the survey and answered more than page two of seven, the surveys were included in the study.
-Therefore, 34 complete and partially completed surveys were used in the analysis.
-Figure 4.
-Distribution of responses from online survey participants.
-Q1 to Q3 were answered by participants with disabilities, Q4 to Q6 by family and friends, and Q7 and Q8 from healthcare professionals.
-[…]
+Of the 71 participants, 36 surveys with no responses (N/A) had to be excluded.
+19 participants completed the survey, and 15 partially completed surveys were included when they answered at least Q1 to Q12 (referring to the question indicators stated in Appendix A), due to their given insights into daily life.
+Therefore, the analysis used 34 complete and partially completed surveys.
+We assume that the large number of empty surveys resulted from people skipping through the survey out of curiosity.
+Additionally, some institutions reported that they conducted an internal ethical evaluation before distributing the survey to patients and staff.
+These surveys were automatically saved, even if they were not finished.
+The group distribution of the completed surveys was 10 PSMD, 3 family members or friends (fam./fr.), and 6 medical professionals (med.
+prof.).
+In comparison, the in the incomplete surveys 4 PSMD, 6 fam./fr., and 5 medical professionals participated.
+The demographics can be found in Table 1, which summarizes the results of Q1 to Q3.
+The general and system-related questions varied by group to identify different viewpoints on the system as shown in Figure 3.
+Table 2 shows the distribution of questions related to the corresponding topic.
+The results of Q4 to Q15 will be presented in the following topics.
+4.1.1. Support situation
+For this topic, questions were asked to evaluate the participants’ personal situations and whether there were any obstacles to filling out the survey.
+Only two of the PSMD group received assistance filling out the survey (Q4(PSMD)).
+The reason was not stated.
+Due to limitations in hand and upper body movement, and a lack of access to suitable assistive devices, it can be assumed that the assistance was necessary (Q8, Q9).
+Relatives and friends only stated once that the person they cared for filled out the survey.
+Therefore, the risk of bias due to conferring between participants is low.
+4.1.2. Level of disability and assistive devices
+Q6 revealed that two participants experienced variability in their disability-related limitations from day to day, underscoring the importance of considering individuals with episodic disabilities.
+These individuals may particularly benefit from multimodal control approaches that can adapt to varying abilities, as discussed by Bhattacharjee et al.
+and Nanavati et al.
+(Bhattacharjee et al., 2020; Nanavati et al., 2024b).
+Regarding Q8, participants differed greatly in their motor abilities and, in most cases, their level of disability remained unchanged (Q7: 12 out of 15).
+In the other three cases, the level of disability improved, which could indicate stroke rehabilitation.
+This emphasizes that, for this group, the presented control method may only be applicable for a short time before switching to other control inputs, such as a joystick.
+This emphasizes the importance of systems that are easy to learn and apply, assisting in short recovery times.
+Relatives and friends stated that three participants had access to assistive devices, in the form of a patient lift and a repositioning aid.
+4.1.3. Robotic arm use
+Of all the participants asked, only one stated that she had already used a robotic arm controlled by a chin switch.
+In the context of Q7(PSMD) and Q8(PSMD), it may be that the level of disability of the other participants does not qualify them to receive a robotic arm from healthcare providers, or they may receive assistance from caregivers or family, which is more common in Germany, Switzerland, and Austria.
+Approval for a medical device by healthcare providers in Germany can take several months or, in rare cases, years.
+This indicates that the use of robots is uncommon and can greatly bias the participants’ perception of the system due to its novelty.
+4.1.4. WoZ satisfaction
+After viewing the video, the participants were asked if they or the person they cared for would be able to use the system.
+More than half of the participants confirmed this (17 out of 34; Q13 - PSMD, Q14 - Rel./Fr.
+and Med.
+Prof.).
+Only three participants stated the opposite.
+No reasons were given.
+Medical professionals and family members were also asked if they were satisfied with the setup process (Q11 - Rel./Fr.
+and Med.
+Prof.).
+Five stated they were satisfied, while three were not.
+One reason given was the amount of time it took to set up.
+It might also be influenced by the desire for more information about the system.
+As Figure 4 shows, the high number of N/A answers in the partially completed surveys occurred after the video.
+This may be because participants were unable to return to the website after watching the video.
+Due to missing feedback, this is only a presumption.
+Figure 4. Distribution of responses from online survey participants. Q4, Q9 and Q13 were answered by PSMD, Q6, Q11 and Q14 by family and friends and medical professionals.
 This bar chart shows the distribution of responses of online survey respondents.
 Eight questions are presented that highlight the most interesting findings.
-People with disabilities were asked the following three questions: First question: Did you have assistance in completing the survey?
+PSMD were asked the following three questions: First question: Did you have assistance in completing the survey?
 Responses: fully completed surveys 2 yes, 12 no, 0 not answered.
@@
 Responses: fully completed surveys 2 yes, 1 no, 6 not answered.
-Caregivers were asked the following two questions: First question: Are you satisfied with the setup of the system?
+Medical professionals were asked the following two questions: First question: Are you satisfied with the setup of the system?
 Responses: fully completed surveys 4 yes, 1 no, 6 not answered.
@@
 Responses: fully completed surveys 6 yes, 0 no, 5 not answered.
-Figure 4 summarizes the results of the survey.
-14 participants identified themselves in the group of people with disabilities, eight in family members, and 11 in caregivers.
-Only once did the respondent not select one of these target groups (”other”).
-The authors decided to include the participant in the family group because of the affiliation description ”Family but without care share”.
-The results show that most people with disabilities were able to complete the survey on their own (Q1), leading to the conclusion that they have an assistive device that allows them to use a smart device.
-This conclusion is based on the participants’ statement that most of them cannot move their hands (10 out of 14).
-After viewing the video, participants were asked if they or the people they care for would be able to use the system.
-More than half of the participants confirmed this (17 out of 32; Q3, Q6, and Q8).
-Only three participants stated the opposite.
-The caregivers and family members were also asked if they were satisfied with the setup process (Q5 and Q7).
-Five stated they were satisfied, while three were not.
-One reason given was the amount of time it took to set up.
-Participants were also asked what tasks they would like to accomplish with the system.
-The group of people with disabilities gave a total of 14 responses: Eating (3 mentions), drinking (2 mentions), reaching for cutlery, pushing buttons (elevator or electric door), applying makeup or brushing their hair, cleaning, scratching, opening a bottle, and lifting objects of various weights from the floor (2 mentions).
-Family members and caregivers were asked this question twice.
-First, they were asked to identify situations in which a robotic system could assist with everyday tasks while they were unaware of the system demonstration.
+4.1.5. Task preference
+Participants were asked what tasks they would prioritize to accomplish with the system.
+The group of PSMD gave a total of 14 responses: Eating (3 mentions), drinking (2 mentions), reaching for cutlery, pushing buttons (elevator or electric door), applying makeup or brushing their hair, cleaning, scratching, opening a bottle, and lifting objects of various weights from the floor (2 mentions).
+Family members and medical professionals were asked this question twice.
+First, they were asked to identify situations in which a robotic system could assist with daily tasks, while they were unaware of the system demonstration.
 Second, they were asked about the presented system.
-There were 13 responses to the first question.
-They rated the following tasks as useful: Making the bed, brushing teeth, assisting with eating, thromboses prevention, putting on socks, documentation, going to the bathroom, doing laundry, cleaning (2 mentions), assisting with walking, assisting with doctor’s appointments and correspondence, silent alarm for service calls, and making phone calls.
-There were 14 responses to the second question: Eating (5 mentions), drinking, turning on lights, operating the aforementioned lift system, brushing teeth, preparing a drink, opening doors, lifting objects (2 mentions), making phone calls and moving objects.
-In comparison both questions show a high potential for robotic systems and digitization.
-In the first case, the focus is on assistance with household tasks and communication, while in the second, user-centered tasks are preferred.
+Thirteen people responded to the first question.
+They rated the following tasks as useful: Making the bed, brushing teeth, assisting with eating, thrombosis prevention, putting on socks, documentation, going to the bathroom, doing laundry, cleaning (2 mentions), assisting with walking, assisting with doctor’s appointments and correspondence, silent alarm for service calls, and making phone calls.
+Fourteen responses were given to the second question.
+Eating (5 mentions), drinking, turning on lights, operating the aforementioned lift system, brushing teeth, preparing a drink, opening doors, lifting objects (2 mentions), making phone calls, and moving objects.
+Both questions demonstrate a high potential for robotic systems and digitization.
+The first focuses on assistance with household tasks and communication, while the second focuses on user-centered tasks.
+Figure 5 shows the distribution of task preferences mentioned by the participants.
+Figure 5. Classification of all mentioned tasks in relation to the presented system. 28 tasks were classified into six categories containing the tasks mentioned by the participants.
+This pie chart shows the classification of all mentioned tasks in relation to the presented system.
+28 tasks were grouped into six categories containing the tasks mentioned by the participants and are shown in percentage.
+The largest category is nutrition with 52%, the second largest category is daily tasks with 19%, the third is mobility with 11%, the fourth is two categories with 7% each described by the terms personal hygiene and well-being.
+Finally, tasks classified as communication are mentioned with 4%.
 4.2. Hands-on study
 4.2.1. Gaze behavior analysis
-The analyses of dwell times in relation to object selection events was used to find technical requirements on the eye-tracking controller.
-Assumptions such as correlations between object size and dwell time duration were investigated for the dwell time trigger anticipated in the design.
-Table 1 shows the distribution of participants’ choice of object selection and information about dwell time.
+Analyses of dwell times in relation to object selection events were used to determine the technical requirements for the eye-tracking controller.
+Assumptions were investigated, such as the correlation between object size and dwell time duration.
+Table 3 shows the distribution of participants’ object selection choices and dwell time information.
 A visual attention map is shown in Figure 6.
-The chosen interval was 1043 ms before the robot started to move and was based on the time steps made by the evaluation software.
-Both the data in Table 1 and in Figure 5 used the set intervals.
-In this time period, events such as searching for the object or deciding which object to select were excluded.
-It was ensured that the participant only focused on the selected object.
-The average duration of the test was 2 min 41 sec (SD = 9.8 sec).
-The average dwell time was calculated to determine the maximum dwell time limit for the controller, including the time from fixation start until the robot moved.
-The longest possible dwell time for the participant’s gaze in this setting was found to be 5.612 sec.
-In this interval, the gaze did not leave the AoI.
-It indicates that dwell time triggers can be set to high durations, but in the following limitations of the maximum dwell time are presented.
-In this study, we assumed that the dwell time is longer the larger the object, due to the larger AoI.
-Accuracy errors would have less weight than in the small AoIs.
-The correlation between object sizes and corresponding dwell times were measured and are given in Figure 6.
-The bottle has a higher standard deviation in dwell time than the cup, implying longer dwell times in some participants’ trials.
-As seen in Table 1 this is also visible in the dwell times above 500 ms.
-Reasons why the correlation may not apply are, for example, the fixation of the bottle cap, which is closer to the edge of the AoI than the cup handle.
-The gaze is more likely to leave the AoI, resulting in lower dwell times and indicating the influence of intention.
-In comparison, the fork and the block have approximately the same average dwell time and SD, despite having different object sizes.
-Our first hypothesis was extended that the geometry of the object also influences the dwell time.
-This influence is indicated by the small width of the fork handle, being less than 1 cm.
-When accuracy errors occur, it is more likely that gaze is tracked outside the AoI.
+The chosen interval was 1,043 ms before the robot started moving, based on the time steps made by the evaluation software.
+Both data in Table 3 and in Figure 6 represent the intervals.
+During this time period, events such as searching for the object or deciding which object to select were excluded.
+Participants were instructed to focus solely on the selected object, to investigate whether natural gaze behavior exhibits gaze cues that hinder implementation or result in insufficient accuracy.
+The average duration of the test was 2 min, 41 sec (SD = 9.8 sec).
+The average dwell time was calculated to determine the maximum limit for the controller’s dwell time, including the time from the start of fixation until the robot moved.
+In this setting, the longest possible dwell time for the participant’s gaze was found to be 5.612 seconds.
+During this time, the gaze did not leave the area of interest (AoI).
+This indicates that dwell time triggers can be set to long durations.
+However, the following limitations show that such long dwell times are rare.
+In this study, we assumed that dwell time would be longer for larger objects due to larger AoIs.
+Accuracy errors would have less impact on AoI interactions than in the small AoIs.
+The correlation between object sizes and corresponding dwell times was measured and is given in Figure 7.
+The dwell time of the bottle has a higher standard deviation than the cup, implying more occurrences of longer dwell times.
+As seen in Table 3, this is also visible in the dwell times above 500 ms.
+Reasons why the correlation may not apply are the fixation of the bottle cap, which is closer to the edge of the AoI than the cup handle.
+The gaze is more likely to leave the AoI, resulting in lower dwell times, which indicates the influence of user intention.
+In comparison, the fork and the block have similar average dwell times and standard deviations, despite their different sizes.
+We extended our initial hypothesis to include the influence of geometry of the object on the dwell time.
+We tested this by evaluating missed hits and gaze behavior with a fork handle that was less than 1 cm wide.
+When accuracy errors occur, it is more likely that gaze is tracked outside the fork’s AoI.
 Table 3. Information about distribution of selection by participants, dwell times and missed AoI hits.
@@
 […]
-The figure shows the experimental setup shown in Figure 1.
-It is overlaid with a heat map visualizing the attention of all participants during task selection (absolute duration).
-The heat map is visible on the objects and, in some cases, the area around the objects.
-Task 1: Three hotspots of low and medium intensity are shown on the light switch.
-There are few fixations outside the light switch towards the center of the experimental setup.
-Task 2 - Pick and place: The block shows a high intensity hotspot at the top of the block in the left corner toward the center of the experimental setup.
-The hotspot extends beyond the boundary of the object.
-A few fixations to the right of the block are visible in the vicinity of the block.
-The platform shows an elliptical hotspot perpendicular to the height (y-axis) of the figure with low, medium, and high intensity.
-Few fixations were found outside the platform above the object.
-Task 3 - Filling a cup: The cup’s high-intensity hotspot is centered on the top part of the handle and extends slightly beyond the object’s outline.
-The bottle has a vertically elongated hotspot that starts at a third of the height of the bottle from the bottom and continues to the bottle cap, slightly exceeding the objects outlines.
-Although only low intensities are shown, the highest intensity is at the bottle cap.
-Task 4 - Picking a food item: The hotspot on the food item is off-center from the food item itself and above the food item.
-The hotspot is about the same size as the food item and covers approximately half of the food item itself.
-A second low-intensity hotspot is shown to the right of the food item, between the food item and the fork.
-The fork has two hotspots at the base of the fork tines and slightly below in 1/6 of the fork handle.
-Both hotspots extend across the width of the fork.
 Figure 7.
@@
 The line in the box plots marks the median, confidence intervals are given at 95%.
-[…]
 This graph shows the correlation between object size, missed hits and dwell time for the selectable objects.
@@
 Line plot results for missed hits of the AoI of each object: Bottle: 0, cup: 0, fork: 5, food item: 7, platform: 1, block: 5, light switch: 4.
-The influence of gaze direction was found by examining the low dwell times of the food item placed closest to the user.
-Looking at the gaze distribution in relation to the scene camera coordinate system, as shown in Figure 7, it is evident that participants tended to move their head to center their gaze when looking at a peripheral object to the right or left.
-In contrast, participants tended not to move their head when looking down.
-The detected gaze location was less accurate, due to higher scatter closer to the eye-tracker’s periphery.
-Physiological impact cannot be excluded such as shadowed pupils by the eyelashes.
+Examining the low dwell times of the food item placed closest to the user revealed the influence of gaze direction on the dwell time.
+Looking at the gaze distribution in relation to the scene camera’s coordinate system reveals that participants tended to move their heads to center their gaze when looking at objects to the right or left of them, as shown in Figure 8.
+In contrast, participants tended not to move their heads when looking down.
+The detected gaze location was less accurate due to the greater scatter closer to the eye-tracker’s periphery.
+Physiological factors, such as shadowed pupils by the eyelashes, cannot be excluded to impact dwell time.
 This resulted in seven missed AoI hits for the food item.
-Significant correlation was investigated between dwell time, object size, and gaze direction towards the object.
+Significant correlations were investigated between dwell time, object size, and gaze direction towards the object.
 The Shapiro-Wilk test (H = 0.771, df = 97, p = ¡0.001) and the Levene test (H = 2.658, df1 = 6, df2 = 90, p = 0.02) showed no normal distribution or homogeneity.
@@
 Spearman’s $\rho$ was used to analyze the correlation between object size, object position, and dwell time.
-The frame in Figure 7 was evenly divided into five-by-five sectors in horizontal and vertical directions, so that gaze location could be described by a coordinate map.
-The analysis showed that dwell time and object size ($\rho=0.373$, $p<0.001$) as well as dwell time and offset in y-direction ($\rho=0.263$, $p=0.009$) were positively correlated.
+In Figure 8, gaze location was evaluated by applying a coordinate map visualized by the equally distributed sectors.
+The analysis showed that dwell time and object size ($\rho=0.373$, $p<0.001$) were positively correlated, as well as dwell time and offset in y-direction ($\rho=0.263$, $p=0.009$).
 The correlation between the offset values in y-direction and x-direction is weaker and does not correlate with dwell time ($\rho=0.160$, $p=0.116$).
@@
 Gaze direction was mostly centered and downwards while task selection.
-[…]
 This figure shows the heat map visualization of gaze position in the field of view of the eye-tracking camera during task selection.
@@
 4.2.2. Perception of hands-on application
-In general, satisfaction with the system was expressed in the hands-on study.
-Participants responded with following comments: 9 positive, 3 negative, 9 times positive and negative, and 3 times with no comments.
-All statements can be found in Appendix A.
-Differences were found between the responses of people who work regularly with robotic arms (called prior-experience group below) and people who have no experience with robotic arms (called non-prior-experience group).
-The non-prior-experience group gave mostly positive comments and focused more on the description of tasks, while the prior-experience group focused on robot kinematics, robot behavior, and user safety, which partially included negative aspects.
-While the non-expert group reported being interested, entertained, satisfied, and impressed, they also perceived confusion and uncertainty.
-The task design itself caused confusion, as participants were not sure how to react to the glass presented by the robot.
-Considering a real-life application of filling the glass, one of the participants expressed fear of overfilling the glass.
-Regarding the robotic system, they were satisfied with good reaction time and intuitive control.
-The velocity was ambivalent, with some participants saying it was just right and others saying it was too slow.
+In general, the hands-on study revealed satisfaction with the system.
+In the unstructured interview, the participants’ comments were evaluated based on whether they were positive or negative regarding the robot’s performance.
+The results were as follows: nine positive comments, three negative comments, nine comments that were both positive and negative, and three comments that were neither positive nor negative.
+All statements can be found in Appendix B.
+Differences were found between the responses of people who regularly work with robotic arms (the prior-experience group) and those who have no experience with robotic arms and robots in general (the non-prior-experience group).
+The non-prior-experience group gave mostly positive comments and focused on describing the tasks, while the prior-experience group focused on robot kinematics, robot behavior, and user safety.
+The non-expert group reported feeling interested, entertained, satisfied, and impressed.
+At the same time, they reported feeling confused and uncertain.
+The task design itself caused confusion because participants were unsure how to react to the glass presented by the robot.
+Considering a real-life application of pouring a drink, one participant expressed fear of overfilling the glass.
+Regarding the robotic system, participants were satisfied with good reaction time and intuitive control.
+Negative comments concerned the velocity, which created ambivalent reactions.
+Some participants stated the velocity was just right, and others saying it was too slow.
 Importantly, users reported feeling safer because the robot’s movements were limited to the edge of the table.
-The fork movement may have had an unsettling effect on the person, as they reported that the fork was not close enough to eat or at the right height for their mouth.
-The prior-experience group comments were mostly negative about the robot’s trajectory and the measurement setup.
-For example, the gripper could crush someone’s hand if the user reached into the setup, due to the lack of pressure sensors.
-The placement of the emergency stop was a concern for the well-being of future users due to its distance from the participant.
-The shaky movements of the robot, the non-optimized trajectories, and in one case the robot accidentally touching the bottle while moving to the next object were mentioned.
+The fork movement toward the participant had an unsettling effect, as they reported that the fork was not close enough to eat, or at the right height for their mouth.
+Comments from the prior-experience group were mostly negative about the robot’s trajectory and the measurement setup.
+For instance, the gripper could potentially crush someone’s hand if the user reached into the setup due to the lack of pressure sensors.
+The placement of the emergency stop was a concern for the well-being of future users, due to its distance from the participant.
+They also mentioned the robot’s shaky movements, non-optimized trajectories, and one instance where the robot accidentally touched the bottle while moving to the next object.
 The perceived motor vibration was uncomfortable and the sound of the dropped fork was too loud.
-Reaction time and object recognition were mentioned as positive.
-In addition, two participants from both groups stated that they felt uncertain that the robot would stop the task if they looked away or at other objects while the robot was performing the task.
+Reaction time and object recognition were mentioned as positive, implying that the Wizard of Oz design was convincing.
+Additionally, two participants from both groups stated that they were uncertain whether the robot would stop the task if they looked away or at other objects while the robot was performing it, a concern that must be addressed in the control design.
 5. Discussion
-In Section 5.1 two of three usability questions and in Section 5.2 two of the three functionality questions are answered asked in the Methods.
-In Section 5.3 the last question of both studies will be discussed, which combines the perception of the online study and hands-on study.
-The resulting adaption features to improve the initial design are given in Section 5.4.
+In Section 5.1, two of the three usability questions and in Section 5.2, two of the three functionality questions are answered.
+In Section 5.3, the last question of both studies will be discussed, which combines the perception of the online study and hands-on study.
+Feature improvements to the initial design are presented.
 5.1. Usability
 5.1.1. Are potential users able to use the system?
-Yes, the survey showed in Q3, Q6, and Q8 (Figure 4) that most participants are able to use the system based on the video presentation (Answers: 17 yes, 3 no, 14 N/A).
-The high amount of N/A results were a result of including partially filled out surveys.
-In the open-ended questions, statements were made that the participants would like to have more information about the system to better evaluate the maintenance needed and the function of the system.
-These requests did not imply that the participant would change his or her mind after receiving the additional information.
-One participant stated “I wonder if the system can be used by people with intellectual disabilities due to their reduced cognition”.
-This underlines the importance of considering a variety of disabilities in the design process of such systems in order to make them accessible to a wide range of users.
-It highlights again the necessity to involve the community in the design process.
-5.1.2. Which tasks are preferred so that the robot has the greatest benefit in everyday life?
-The categories of eating, drinking, environmental interactions, pick and place, personal hygiene, and miscellaneous tasks were found in the literature and are listed in Section 2.2.
-The participants’ responses from Section 4.1 were sorted into these categories that lead the distribution in Figure 8.
-The most requested task in the online survey that the robot should perform is assisting to eat or drink, followed by environmental interaction, which includes turning lights on and off, pressing buttons, opening doors, and pick and place objects.
-Picking up and moving objects was mentioned in different contexts such as moving heavy objects out of the way.
-Personal hygiene consisted of a variety of different tasks including doing makeup and hair, brushing teeth, and cleaning.
+Yes, the survey showed in Q13(PSMD) and Q14(Fam./Fr.
+and Med.
+prof.) (Figure 4) that most of the participants are able to use the system based on the video presentation (Answers: 17 yes, 3 no, 14 N/A).
+The high amount of N/A results were the result of including partially filled out surveys.
+The limiting factors that prevent the system from being used every day were not stated.
+In the open-ended questions, participants stated that they would like more information.
+They also said that they would like more information about the system to better evaluate its maintenance and function.
+These requests did not imply that the participants would change their minds after receiving the additional information.
+One participant stated, “I wonder if the system can be used by people with intellectual disabilities due to their reduced cognition”.
+This highlights the importance of considering various impacts of disabilities during the design process to make such systems accessible to a wide range of users.
+It again highlights the necessity of involving the community in the design process.
+Deeper issues, such as how the control works and whether wrong selections can occur due to natural gaze behavior, were not questioned by the participants, contrasting the results from the hands-on study.
+Section 5.3 shows participants’ concerns about interacting directly with the system, highlighting the benefit of dual investigations in robot perception.
+5.1.2. Which tasks should be prioritized to maximize the robot’s usefulness in everyday life?
+The results, shown in Figure 5, the most requested tasks are assisting while eating or drinking, followed by environmental interaction, which includes turning lights on and off, pressing buttons, opening doors, and picking and placing objects.
+Personal hygiene consisted of a variety of different tasks, including doing makeup and hair, brushing teeth, and cleaning.
 Miscellaneous tasks included making phone calls and scratching themselves.
-These results can be corroborated with other findings (Chang and Park, 2003; Stanger et al., 1994; WHO, 2010).
-The variety of tasks expressed shows that participants were not biased by the tasks shown in the video.
-Figure 8. Classification of all mentioned tasks in relation to the presented system. 28 tasks were classified into six categories containing the tasks mentioned by the participants.
-[…]
-This pie chart shows the classification of all mentioned tasks in relation to the presented system.
-28 tasks were grouped into six categories containing the tasks mentioned by the participants and are shown in percentage.
-The largest category is nutrition with 52%, the second largest category is everyday tasks with 19%, the third is mobility with 11%, the fourth is two categories with 7% each described by the terms personal hygiene and well-being.
-Finally, tasks classified as communication are mentioned with 4%.
-Brushing one’s hair, applying makeup, or scratching are everyday tasks that require close human-robot interaction while ensuring the safety of the user.
+Brushing one’s hair, applying makeup, or scratching are daily tasks that require close human-robot interaction while ensuring the safety of the user.
 The shared control approach presented would require face and body recognition if the user is unable to move their head to a certain position.
@@
 The aim must be to present this information in simple language so that people with physical and cognitive disabilities can understand it and make decisions.
+The prioritized tasks and the tasks realized from the literature included in Section 2.3 were summarized.
+A comparison of the results with the findings in the literature shows partial agreement.
+Table 4 compares findings from the literature with our results.
+In the ”Prioritization” column, we summarized the works of Beer et al., Nanavati et al., Petrich et al., Smarr et al., Stanger et al., Chang et al., and Jardón et al.
+(Beer and Akanksha Prakash, 2011; Petrich et al., 2022; Nanavati et al., 2024b; Smarr et al., 2012; Stanger et al., 1994; Chang and Park, 2003; Jardón et al., 2012).
+These works present or review task prioritization in activities of daily living (ADLs), involving either elderly people or people with disabilities.
+It is important to acknowledge that these works include varying participant groups, as well as varying robot types.
+To the best of the authors’ knowledge, no review or work solely focuses on PSMD and robotic arms.
+The ”Integration” column summarizes reviews and individual studies concerning the execution of tasks with assistive robots.
+It summarizes the outcomes of the following works, which were also presented in Section 2.3 (Nanavati et al., 2024b; Hagengruber et al., 2025; Dragan and Srinivasa, 2013; Huang and Mutlu, 2016; Zhang et al., 2025; Bien et al., 2003; Canal et al., 2021; Bhattacharjee et al., 2020).
+Some tasks were excluded from the literature when they could not be performed with a robotic arm (e.g., getting around or into bed, going out, toileting, managing money, leisure and recreation, and shopping), as well as outdated tasks (e.g., removing paper from printers and fax machines).
+Table 4.
+Overview of tasks mentioned by participants and representation within literature.
+Tasks mentioned by users are presented to their according categories in Figure 5.
+Fit describes if the task can be accomplished without additional assistive devices to a robotic arm (triangle shows partial accomplishment).
+The comparison is done between our results, task prioritization found in other works and integrated tasks in existing works.
+The table presents an overview of daily-life task categories, specific tasks within each category, and how these tasks are represented in the collected user responses and in prior literature on assistive robotic arms.
+Categories include Eating, Drinking, Environmental Interaction, Personal Hygiene, and Miscellaneous Tasks.
+For each task, the table indicates whether it can be performed by a robotic arm without additional assistive devices, and compares three aspects: how frequently the task was mentioned by participants in the hands-on study, how strongly participants in related literature studies preferred the task, and how strong the research interest in realizing the task in existing assistive robotic systems is.
+Eating and drinking are both considered feasible with a robotic arm alone and are frequently mentioned by users, commonly prioritized in literature, and often integrated in existing systems.
+Environmental interaction tasks include preparing meals, using switches and buttons, opening doors and drawers, pick-and-place actions, and carrying objects.
+All of these tasks can be accomplished with a robotic arm alone.
+All tasks except preparing meals are mentioned by participants in this study.
+They are highly prioritized by participants in other studies, and commonly integrated in existing systems, with Switches, Buttons and carrying objects showing slightly less frequent integration.
+Personal hygiene tasks include makeup or shaving, combing hair, brushing teeth, dressing, and bathing or grooming.
+All are feasible with a robotic arm alone, except bathing and grooming.
+Bathing and grooming are only partially feasible without additional assistive devices.
+They mentioned by participants in this study except Dressing and Bathing which also represents the results from other literature.
+They are less commonly integrated into existing systems.
+Miscellaneous tasks include using phones or electronics, taking medicine, doing household chores, and scratching.
+Household chores are feasible, mentioned in this study and literature and frequently integrated in existing systems.
+Using electronics and scratching are only partially feasible without additional devices.
+Taking medicine is feasible but less frequently mentioned and less often integrated in existing systems.
+Category Tasks Fit Our results Prioritization Integration
+Eating Eating $\checkmark$ $\nearrow$ $\rightarrow$ $\uparrow$
+Drinking Drinking $\checkmark$ $\nearrow$ $\nearrow$ $\uparrow$
+Environmental Preparing meals $\checkmark$ $\uparrow$ $\uparrow$
+Interaction Switches, Buttons $\checkmark$ $\nearrow$ $\uparrow$ $\rightarrow$
+Doors, drawers $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
+Pick and place $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
+Carrying objects $\checkmark$ $\nearrow$ $\uparrow$ $\nearrow$
+Personal Makeup, shaving $\checkmark$ $\nearrow$ $\uparrow$ $\rightarrow$
+hygiene Combing hair $\checkmark$ $\nearrow$ $\nearrow$ $\searrow$
+Brushing teeth $\checkmark$ $\nearrow$ $\nearrow$ $\searrow$
+Dressing $\checkmark$ $\rightarrow$ $\rightarrow$
+Bathing, Grooming $\triangle$ $\rightarrow$ $\nearrow$
+Miscellaneous Using phone, electronics $\triangle$ $\nearrow$ $\rightarrow$ $\rightarrow$
+tasks Taking medicine $\checkmark$ $\rightarrow$ $\searrow$
+House choirs $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
+Scratching $\triangle$ $\nearrow$ $\nearrow$ $\searrow$
+Tasks such as preparing meals, dressing, bathing, grooming, and taking medicine were not mentioned by the participants.
+This may be because these tasks are comparatively less important to end users, as found in the literature.
+In contrast, preparing meals is highly prioritized and researched in other studies.
+Depending on their hobbies or the type of personal assistance they receive, this task may not have been favored by the small group of participants.
+The insight into the eating task is interesting.
+While most participants stated that they strongly preferred to accomplish this task independently, Nanavati et al.
+found that, compared to other ADLs and IADLs, this task was the least prioritized (Nanavati et al., 2024b).
+However, other sources, such as Petrich et al.
+and Stanger et al., found it to be highly prioritized (Stanger et al., 1994; Petrich et al., 2022).
+This discrepancy may be due to the fact that Nanavati et al.’s work reviewed different populations, including people with visual impairments and the elderly, who may be able to eat without assistance.
+Comparing works that address these tasks reveals gaps in interacting with switches and buttons, personal hygiene, scratching, and using electronics.
+Concerns about user safety and privacy regulations might influence the low number of works related to personal hygiene and scratching.
+In contrast, using electronics or phones is becoming more convenient with smart home applications and hands-free technology.
+In conclusion, the chosen task for the presentation aligns well with the literature.
+More focus should be placed on assisting people with personal hygiene tasks.
+The variety of tasks performed demonstrates that the participants were not influenced by the tasks shown in the video.
 5.2. Functionality
@@
 Initial testing showed that 70% of object selections would result in selection success when the dwell time trigger is set to 500 ms.
-Although, dwell times can be shortened to enhance selection success as shown in other literature (Holmqvist and Andersson, 2017), this could lead to an unintentional object selection as described in Section 2.3.
+Although dwell times can be shortened to enhance selection success, as shown in other literature (Holmqvist and Andersson, 2017), this could lead to an unintentional object selection, as described in Section 2.1.
 An appropriate threshold must be explored with the prototype system.
-Accuracy issues are one of the main reasons why gaze was tracked outside of an AoI. Therefore, influences on AoI misses and dwell time were investigated in Section 4.2. Four dependencies were found:
+Accuracy issues are one of the main reasons why gaze was tracked outside of an AoI.
+Therefore, the influences on AoI misses and dwell time were investigated in Section 4.2.
+Four dependencies were found:
   • Object size: As shown in Figure 7, object size correlates significantly with dwell time.
-Small objects like keys are part of everyday activities.
-Excluding such objects from interaction, would lead to major drawbacks in participation in daily life.
+Small objects, like keys, are part of everyday activities.
+Excluding such objects from interaction would lead to major drawbacks in participation in daily life.
 The available workspace is limited by the length of the robotic arm.
 This also limits the relative size of the object in the camera scene.
-With this limit, the resulting accuracy error can be calculated and checked with the desired interactable object size to estimate if this would cause problems.
+With this limit, the resulting accuracy error can be calculated and checked with the desired interactive object size to estimate if this would cause problems.
 In this case, the accuracy of the eye-tracker is $\Delta_{deg}=0.6$°.
@@
 Since all selected objects were larger than 9.34 mm, it is assumed that the prototype would work.
-  • Geometry: The AoI of the fork could lead in a failed selection due to its width.
-Currently, there are two approaches in planning for estimating object outlines.
+  • Geometry: The AoI of the fork could lead to a failed selection due to its width.
+Currently, there are two approaches in planning for estimating object outline.
 Object recognition can be used to create bounding boxes, solving the issue due to the new geometry.
 Overlapping of bounding boxes may occur more often, leading to unintentional object selection and context misinterpretation.
-Object segmentation can lead to boundaries which fit to the object’s outlines.
-In this case the geometry of the object would be traced with high accuracy, solving the bounding box issue, leading back to the initial width issue.
-  • Gaze location towards the object: In the eating task example, participants tend to move their head to the left and right instead of up and down (see Figure 7).
+Object segmentation can lead to boundaries which fit the object’s outline.
+In this case, the geometry of the object would be traced with high accuracy, solving the bounding box issue, leading back to the initial width issue.
+  • Gaze location towards the object: In the eating task example, participants tend to move their heads to the left and right instead of up and down (see Figure 8).
 The resulting higher accuracy error was due to the glasses periphery.
@@
 In the case of the bottle, the task was to fill the glass.
-The bottle needs to be opened first, which explains why the bottle cap was fixated, indicating the hand-eye coordination and intention behind the task as discussed in Section 2.1.
-5.2.2. What are the technical difficulties with the design?
+The bottle needs to be opened first, which explains why the bottle cap was fixated, indicating the hand-eye coordination and intention behind the task, as discussed in Section 2.1.
+5.2.2. What technical difficulties arise from the design?
 Ambiguous object selection can occur due to object overlap and was found in the case of the fork (see Figure 6).
 In addition to the accuracy challenges, robust object selection may be the most difficult challenge in implementing the prototype.
-Task validation methods and emergency stops must be ensured, as found in Section 2.1.
-Approaches such as looking into the corner of the field of view could be challenging to interpret as an emergency stop because it is processed as the same data type as prolonged blinking in the eye-tracking software.
-Another approach describes the inclusion of gaze gestures (unique sequence of eye movements (Alsharif, 2017)).
-It needs to be verified whether such a cue would be able to perform for a person in an emergency situation.
-For task validation, the system needs to be enhanced with feedback methods to communicate the expected task by the robot.
+Other researchers have addressed this issue using 3D gaze tracking to minimize localization errors in approaches involving eye-tracking glasses (Cao et al., 2025; Yang et al., 2021).
+As discussed in Section 2.1, user intent may lead to more robust selection methods by involving multiple objects to estimate task intent.
+These objects may vary in geometry and size, which reduces the impact of incorrect fixations.
+Regarding current approaches using object detection, overlap in bounding boxes of closely arranged objects may lead to larger issues than those found in the area of interest (AoI) evaluation.
+The AoIs were fitted to the outlines of the objects, similar to object segmentation.
+However, bounding boxes have larger areas, which leads to ambiguity when they overlap.
+Section 2.2 presents technical challenges, level of autonomy, and user preferences as factors influencing user acceptance.
+Within these categories, we found indications from the survey and hands-on study of technical difficulties.
+After implementing the suggested control method, the tasks will be selectable as planned.
+However, as Hagengruber et al.
+(Hagengruber et al., 2025) point out, accomplishing subsequent tasks is necessary to realize benefits for everyday usage.
+With this in mind, objects could indicate multiple tasks.
+Solutions integrating machine learning methods to identify gaze patterns, as presented by Yang et al., Jain et al., and Belardinelli, can reduce misinterpretation (Yang et al., 2021; Jain and Argall, 2019; Belardinelli, 2024).
+In the WoZ design, the robot was presented with a high level of autonomy.
+As discussed in Section 2.2, it is questionable whether participants will accept the system as intended.
+In the hands-on study, participants vocalized their safety concerns, especially in the group with no prior robot experience.
+Arbitration methods can improve perception.
+This is especially important, since robot perception changes with regular use, and the level of autonomy can be reduced until the user is confident enough to operate the robot with a higher level of autonomy.
+Additionally, settings should be integrated that allow users to adjust the robot’s velocity to their preferred level.
 5.3. Differences between Online Study and Hands-On User Study
 5.3.1. Were the participants satisfied with the presented system?
-In both studies, participants were mostly satisfied with the system. The challenges that were uncovered are summarized in Section 5.4.
-5.3.2. Does participant perception vary greatly between the hands-on study and the online survey?
-The unstructured in-person interview with participants in the hands-on study and the open-ended questions in the online survey, provided more diverse feedback on this shared control design.
-The results are consistent with and extend current literature in this area.
-It adds information for the design of gaze-controlled assistive robotic arms.
-Perceptions of the system do not vary but, highlight different aspects of the design.
-Involving the stakeholder group yielded insights about everyday tasks and their daily life with their disability.
-The prior-experience group revealed potential drawbacks in the technical design process, such as concerns about safety, kinematic selection, and trajectory planning.
-Although the non-prior-experience group did not have in-depth feedback on certain aspects of the life with a disability or technology, they helped build a larger pool of participants to measure gaze behavior.
-In addition, their perception of the robot was more curious and at the same time hesitant than that of the prior-experience group, due to the novelty of robot interaction.
-Since robotic arms are rarely available to people with disabilities, it is reasonable to assume that in a real-world test, people with disabilities would respond in a initial robot interaction similarly to the non-prior-experience group.
-5.4. Adaptation of the initial Shared Control design
-Based on these results, the design can be refined to meet the participants’ expectations of the system.
-The next step is to automate the experimenter’s control task so that the robot can autonomously perform tasks selected by the user.
+In both studies, the participants were mostly satisfied with the system.
+We found in both studies suggestions for improvement.
+The design can be refined to meet the participants’ expectations of the system.
 The framework must have the following features to realize the system:
   • Gaze detection on objects: Due to the outcomes in Section 5.2, several approaches are followed to detect objects in the scene camera of the eye-tracking glasses.
-Unintentional selection and context misinterpretation will be estimated with the prototype.
+The prototype will estimate unintentional selection and context misinterpretation.
   • Task availability: The survey showed that the desired tasks were in line with the participants’ expectations.
 Eating was the most frequently mentioned task.
-However, handling a fork is a safety risk, when the robot manipulates the fork autonomously and has to be handled appropriately.
+However, handling a fork is a safety risk when the robot manipulates the fork autonomously, and has to be handled appropriately.
   • Task selection: The expected prototype will implement objects associated with multiple tasks.
@@
 As mentioned in Section 2.1, feedback strategies need to be evaluated and integrated into the control to ensure the user.
-  • Trajectory, grasping and collision avoidance: The robot has to perform the task by itself.
-In the current version, the robot’s trajectory was programmed and the objects were placed at the same location for each test.
-In a real-world scenario this would not meet the user’s expectations.
-Solutions were found as presented in Section 2.3) and depend on the approach chosen for object gaze detection.
-  • Adaptation: On behalf of the ambiguous results of the robot’s velocity perception, an adaptation setting will be included in the prototype so that users can adapt the system to their liking such as robot velocity.
+Strategies have to be evaluated if user intent methods assist in solving issues selecting between multiple tasks.
+  • Trajectory, grasping, and collision avoidance: The robot has to perform the task by itself.
+In the current version, the robot’s trajectory was programmed, and the objects were placed at the same locations for each test.
+In a real-world scenario, this would not meet the user’s expectations.
+Solutions were found, as presented in Section 2.1), and depend on the approach chosen for object gaze detection.
+  • Adaptation: Due to the ambiguous results of the robot’s velocity perception, settings will be included in the prototype so that users can adapt the system to their liking, such as robot velocity.
+In future work, the experimenter’s control has to be automatized so that the robot can autonomously perform tasks selected by the user.
+5.3.2. Does perception vary between the hands-on study and the online survey?
+The unstructured, in-person interviews with participants in the hands-on study and the open-ended questions in the online survey provided more diverse feedback on this shared control design.
+The results are consistent with, and extend, current literature in this area.
+It adds information for the design of gaze-controlled assistive robotic arms.
+Perceptions of the system do not vary, but highlight different aspects of the design.
+Involving the stakeholder group yielded insights about daily tasks and their daily life with their disability.
+The prior-experience group revealed potential drawbacks in the technical design process, such as concerns about safety, kinematic selection, and trajectory planning.
+Although the non-prior-experience group did not have in-depth feedback on certain aspects of life with a disability or technology, they helped build a larger pool of participants to measure gaze behavior.
+In addition, their perception of the robot was more curious and, at the same time, hesitant than that of the prior-experience group due to the novelty of robot interaction.
+Since robotic arms are rarely available to PSMD, it is reasonable to assume that, in a real-world test, PSMD would respond in an initial robot interaction similarly to the non-prior experience group.
 6. Limitations
-Additional questions: The safety concerns were not reflected in the responses to the online survey.
-Including questions such as ”Would you feel safe interacting with the robot?” would provide a more comprehensive understanding of this topic.
-However, other studies show that potential users are more likely to accept the risks of using the system than to reject it and become dependent on caregivers or family members (Baumeister et al., 2021).
-This can also be seen here, as the robot-experienced group is aware of the risks of human-robot interaction but is not fearful of interacting with the robot.
-Specific Wizard of Oz design: The Wizard of Oz system is an eye-tracking-driven design in which objects are selected without the need for a GUI.
-This approach is uncommon since prototype systems face challenges such as interpreting the desired task, error estimation due to head movements, and a lack of feedback modalities (Fischer-Janzen et al., 2024).
-Researchers using these adaptation suggestions should check whether the findings can be applied to different input designs.
-Generalized approach: This approach was not tailored to specific disabilities, so it may lack details necessary to adapt to individual users’ needs.
-Future work can use this generalization to more easily locate these demands.
+Online study: The qualitative analysis of the online survey provided limited insight into the participants’ everyday lives, meaning the data cannot be generalized to a larger population.
+To counter this issue, we enriched and compared our information with suitable literature.
+For future studies, however, we have decided to conduct interviews with participants to minimize potential hurdles due to a lack of assistive devices or display problems on end devices.
+Q4 was tailored to evaluate whether hurdles exist.
+No indications were found.
+Additionally, no emails stating technical hurdles were received from medical professionals, family, or PSMD.
+The online survey might have benefited from additional questions.
+The safety concerns were not reflected in the survey responses.
+Including questions such as ”Would you feel safe interacting with the robot?” would have provided a more comprehensive understanding of this topic.
+However, other studies show that potential users are more likely to accept the risks of using the system than reject it, becoming dependent on medical professionals or family members (Baumeister et al., 2021).
+This can also be seen here, as the group with robot experience is aware of the risks of human-robot interaction, but is not fearful of interacting with the robot.
+The survey did not ask about the occupation of the medical professionals.
+Therefore, opinions may differ depending on the occupational field (e.g., doctors, caregivers, and nurses).
+The online survey was designed for quantitative analysis.
+Due to the limited number of participants, a qualitative analysis was performed.
+Therefore, not all answers were discussed in detail, since they are not representative (e.g., Q5 rel./fr.: time spent on care).
+These results can be accessed in the supplementary materials.
+As discussed in Section 3.3, participants did not interact directly with the robot, which might affect perception.
+Specific Wizard of Oz design: The Wizard of Oz system does not vary in intent interpretation due to its hard-coded trajectories.
+Variable behavior was excluded from this study, since comparing outcomes between participants would not be applicable.
+However, varying behavior in shared control changes the perception of the system, as shown in (Aronson et al., 2018).
+We presented a design driven by eye-tracking, in which objects are selected without the need for a GUI.
+This approach is uncommon, since prototype systems face challenges such as interpreting the desired task, estimating errors due to head movements, and lacking feedback modalities (Fischer-Janzen et al., 2024).
+If differences in control design exist in relation to the WoZ design, the results may not be transferable to other controls.
+Generalized approach: Since this approach was not tailored to specific disabilities, it may lack the details necessary to adapt to the needs of individual users.
+Future work can use this generalization to more easily identify these requirements.
+Transfer to other eye trackers: Only the Tobii Pro Glasses 3 eye tracker was used, which is more accurate than other eye-tracking glasses on the market.
+However, the effect on the bounding boxes in object detection can vary greatly depending on the model used.
+Due to the scope of this paper, the test was not reproduced with other eye trackers.
 7. Conclusion
-The goal of this work was to obtain feedback on an eye-tracking-driven shared control for Assistive Robotic Arms to identify community expectations, desired tasks, and technical challenges.
-The factors for a successful implementation of a shared control were identified and presented.
-A Wizard of Oz design was created prior to prototyping, as suggested by other researchers.
-Based on this design, two studies were conducted.
-An online survey designed for people with disabilities, their families and friends, and caregivers provided feedback on the community’s expectations for the system, as well as the desired task the robot should be able to perform.
+The goal of this work was to obtain feedback on an eye-tracking-driven shared control for assistive robotic arms in order to identify community expectations, desired tasks, and technical challenges.
+While previous studies offer valuable findings, they vary between robot types and involve alternative stakeholder groups.
+As a result, the available insights are generally only limited transferable and not specific to the control for an assistive robotic arm examined here.
+As suggested by other researchers, a Wizard of Oz design was created prior to prototyping.
+The factors necessary for successfully implementing shared control were identified and presented.
+Based on this WoZ design, two studies were conducted.
+An online survey designed for PSMD, their families and friends, and medical professionals.
+This survey provided feedback on the community’s expectations for the system and the tasks the robot should be able to perform.
 A hands-on study with participants with and without robotics experience provided insight into the technical requirements and feedback differences between the video presentation available for the online survey and the real-world application.
-Gaze data was recorded to measure dwell times and influence dependencies related to the eye-tracking control.
-Dependencies for the design of eye-tracking-driven robot controls were uncovered, ranging from user demands on task prioritization, to accuracy limiting factors in eye-tracking.
-As a consequence, improvements to the initial idea were presented to facilitate the usage of the system.
+Gaze data was recorded to measure dwell times and influence dependencies related to the eye-tracking controls.
+The duality of both studies provided insight into the usability and functionality of the system.
+Task prioritization was contextualized with current literature and approaches that researchers presented to realize shared controls.
+Dependencies were uncovered for the design of eye-tracking-driven robot controls, such as the influence of object size, geometry, gaze location, and user intent on accuracy.
+The impact on object detection or online task intent recognition based approaches were assessed and discussed.
+Consequently, improvements and open challenges related to the initial idea were presented.
 Acknowledgements.
@@
 The study designs were checked to comply to the ACM Publications Policy on Research Involving Human Participants and Subjects.
-The Tobii Pro Glasses 3 was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number INST898/33-1.
+The Tobii Pro Glasses 3 were funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number INST898/33-1.
 Further, we would like to thank Atikkhan Faridkhan Nilgar and Karl Tschurtschenthaler for the great discussions in statistics and eye-tracking evaluation related to this topic.
-We would like to thank Katrin-Misel Ponomarjova to proof read the translations of the questionnaire.
-Lastly, we want to thank Kim Zähringer for her assistance with the robot and video making.
-Appendix A Participant statements of hands-on study
+We would like to thank Katrin-Misel Ponomarjova for proofreading the translations of the survey.
+We want to thank Kim Zähringer for her assistance with the robot and video making.
+Appendix A Overview of Questions asked in the Online Survey
+In the following an Overview of the questions asked to the participants in the online Survey are presented.
+They represent three versions of the survey which would be handed to the individual participant group PSMD, family and friends and medical professionals.
+All related answers can be found in the supplementary materials.
+Table 5. This table presents all questions asked in the online survey. Questions spanning over all three columns indicate questions that were asked to each participant.
+Indicator Questions asked to group
+PSMD Relatives and friends Medical professionals
+Q1 What age group are you in?
+Q2 What gender are you?
+Q3 To which of the groups below do you feel you belong?
+Q4 Do you have assistance from another person in completing this questionnaire?
+Did the person you care for participate in this survey?
+Please estimate the percentage of your patients or clients who have movement difficulties in their arms?
+Q5 If so, what is the reason you need support?
+How much time do you spend on care each day (in hours)?
+Let’s assume that an assistance system could take over all daily tasks.
+Which tasks would you most like to hand over in order to facilitate care?
+Name 1-3 favorites.
+Q6 How often does your impairment affect your daily life? Do you already use assistive devices to help you provide care? -
+Q7 In terms of your physical disability, has your condition improved, worsened, or stayed the same in the last 6 months? If yes, which assistive devices do you use? -
+Q8 Select the body movements that you can control without restrictions - -
+Q9 Do you already use a robotic gripper arm, for example on an electric wheelchair? Does the person(s) you care for already use a robotic gripping arm, e.g. on an electric wheelchair?
+Q10 If so, how do you control it? If so, how does the person control this robotic arm?
+Q11 Are you satisfied with your current system? Are you satisfied with the system in terms of startup and function?
+Q12 If no, indicate why you are not satisfied with the system. If no, why are you not satisfied with the system?
+Q13 In relation to your disability, would it be possible for you to use the system presented in the video every day?
+In which tasks would you or the person you care for most like to be supported by this system in daily care?
+Please name your 1-3 favorites.
+Q14 If no, what changes would you suggest to make the system easier for you to use? Would it be possible for the person you care for to use the system presented in the video every day?
+Q15 In which tasks would you most like to be supported by this system?
+Please name your 1-3 favorites.
+If no, what changes would you suggest to make the system easier to use for you and the person you care for?
+Q16 Do you have any other comments on the questions asked or the system?
+Q17 Would you like to have had more detailed information on how to use the system to answer the questions asked here?
+Q18 If yes, what information would you have liked to have?
+Appendix B Participant statements of hands-on study
 In the following the statements of participants of the hands-on study are listed.
@@
 23 Exp. Robot trembles, trajectory not optimal, bottle badly twisted while handled.
-24 Exp. Fork was recognized well, even though you looked briefly at the bottle during the time you had to aim. The pause between aiming and starting the pick and place task was too long!
+24 Exp. Fork was recognized well, even though you looked briefly at the bottle during the time you had to aim. The pause between aiming and starting the pick and place task was too long.